### v3

# Dynamical Similarity in Multisymplectic Field Theory

Abstract
Symmetry under a particular class of non-strictly canonical transformations may be used to identify, and subsequently excise degrees of freedom which do not contribute to the closure of the algebra of dynamical observables. In this article, we present a mathematical framework which extends this symmetry reduction procedure to classical field theories described within the covariant De-Donder Weyl formalism. Both the Lagrangian and Hamiltonian formalisms are discussed; in the former case, the Lagrangian is introduced as a bundle morphism $\mathcal{L}:J^{1}E\rightarrow\wedge^{m}M$, whilst the Hamiltonian description takes as its multiphase space the restricted multimomentum bundle $J^{1}E^{*}$. The contact reduction process is implemented in a number of particular cases, and it is found that the geometry of the reduced space admits an elegant physical interpretation.
I Introduction
When attempting to formulate a symplectic description of classical field theories, there exist two distinct approaches, both of which have proved highly fruitful in their respective areas of application. On the one hand, the spacetime-covariant formalism, described within the framework of multisymplectic geometry, treats events at a single spacetime point, considering all coordinates (including time) equally [1, 2]; the canonical (or $3+1$) construction, by contrast, aims to study the evolution of a dynamical system from the perspective of the initial-value problem, defined on some suitably-chosen Cauchy surface [3, 4]. In this article, we shall almost exclusively be concerned with the covariant formalism, which in the context of the Hamiltonian description, we refer to as the Hamilton De-Donder Weyl (HDW) formulation [5, 6].
The use of multisymplectic geometry to describe classical field theories presents a number of advantages over its canonical counterpart; in particular, the instantaneous Hamiltonian formalism, almost ubiquitously employed as a precursor to quantum field theory, takes as its canonical variables field functions, evaluated at some fixed instant of time [7]. Consequently, the resulting phase space, upon which we are obliged to work, is infinite-dimensional [8]. Additionally, since the HDW formalism makes no distinction amongst spacetime coordinates when constructing conjugate momenta [9], it is, by its very nature, manifestly covariant [10]. These two features (finite-dimensionality and manifest covariance) have led to significantly hightened research interest within the field of multisymplectic geometry in recent years.
In spite of this, the geometrical formulation is certainly not without its challenges; it is still not fully understood, for example, how the Poisson bracket should be constructed, which, naturally, is an essential prerequisite for most potential geometric quantisation scheme [11, 12, 13]. Some progress has been made in this field [14, 15], and a multisymplectic bracket is well-defined between semi-basic forms on the covariant phase space [16]; however, a fully rigorous construction, applicable to arbitrary forms remains, for the most part, an open question. Additionally, the equal treatment of all spacetime coordinates - which confers the attractive quality of manifest covariance - leads to a discrepancy in the number of field and momentum degrees of freedom: to each field variable, there correspond as many conjugate momenta as dimensions of the underlying spacetime manifold.
Our most precise and accurate description of nature is currently formulated in the language of field theory, both at a classical and quantum level [17]. Indeed, the advent of a field-theoretic viewpoint has arguably been one of the most significant shifts in perspective of the last few centuries. It has been argued by philosophers such as Ismael and van Fraassen that, when seeking a theoretical description of the natural world, one should first construct a framework which amply describes the phenomena of interest, before progressively reducing its superfluous structure, until a minimally-sufficient yet consistent theory remains [18, 19, 20]. It is therefore of interest to examine the structural elements of our field-theoretic models which may be deemed dispensable, and further, to explore any novel insights which are made manifest, as a result of adopting this minimalistic viewpoint. To this end, it is necessary to formulate a precise and rigorous prescription for the identification and subsequent elimination of any structure which is not truly fundamental to our theory.
Our characterisation of mathematical structure as ‘redundant’ shall follow closely the Principle of Essential and Sufficient Autonomy (PESA) [21], which posits that we should seek the minimal ontology that closes the algebra of dynamical observables. This viewpoint is in-line with the ideas of Leibniz, whose Principle of the Identity of Indiscernibles (PII) motivates a theoretical framework, in which only those elements which affect the dynamics of physically-observable quantities are considered [22, 23].
The second facet of our minimalist approach - the elimination of the structure we have labeled as redundant - will rely upon the identification of ‘dynamical similarities’ [24]; these are a class of non-strictly canonical transformations, which leave the dynamics of the observable degrees of freedom invariant. Theories which possess a dynamical similarity may be subjected to a symmetry reduction procedure, in which a single redundant degree of freedom is excised from the ontology. The main objective of the present work is to generalise this process, which is well-established for classical theories of particles, to a field-theoretic context.
This article is structured in the following way: in section (II), we begin by introducing the necessary mathematical framework of multisymplectic geometry, which we apply to the construction of both the Lagrangian and Hamiltonian formulations of classical field theory. We then present a similar treatment of multicontact geometry; together, these constitute the preliminary geometrical ideas required to precisely formulate both conservative and dissipative field theories.
Following this mathematical introduction, we present the concept of dynamical similarity, as defined for theories of particles, constructed on the (co)tangent bundle of some smooth configuration space manifold. As mentioned above, the identification of a dynamical similarity allows a symmetry reduction procedure to be carried out, the details of which are described at the beginning of section (IV). The remainder of this section is then dedicated to presenting a novel generalisation of the reduction procedure for classical field theories, described within the multisymplectic framework.
Dynamical similarities have been studied in a field-theoretic context in [25]; however, this analysis was conducted within the framework of $k$-symplectic geometry [26], which has a number of disadvantages, as compared to the multisymplectic formalism presented here. Of particular note is that, whilst the vast majority of our results focus on regular systems, the framework we construct is a precursor to a full treatment of contact-reduced gauge theories. The fibred manifolds of multisymplectic geometry are the most natural arena in which to develop such a scheme, and thus justifies our choice of formalism. Furthermore, our principal motivation for developing a rigorous method for treating singular theories lies in the study of General Relativity, where it is known that the Einstein-Hilbert action admits a scaling symmetry; we provide further details of this in section (IV.1). Finally, we close with a number of examples, discussing the relevance and practical implications of our results, before providing several open questions and lines of further investigation.
II Multisymplectic Formulation of Classical Field Theories
Our treatment of multisymplectic geometry aims to be self-contained, but by no means exhaustive; since our objective is to describe classical field theories in a covariant manner, much of the material presented throughout this section is adapted with this goal in mind. A more complete and in-depth discussion of the ideas discussed may be found in [27, 28, 29], for example.
II.1 Multisymplectic Manifolds and Multivector Fields
In general, we say that an $m$-dimensional differentiable manifold $\mathcal{M}$ is multisymplectic if it admits a closed, 1-non-degenerate $k$-form $\Omega\in\Omega^{k}(\mathcal{M})$, with $1<k\leqslant m$ [30]. That $\Omega$ is 1-non-degenerate means that for every $p\in\mathcal{M}$ and $X_{p}\in T_{p}\mathcal{M}$
$\iota_{X_{p}}\Omega_{p}=0\quad\quad\implies\quad\quad X_{p}=0$
If $\Omega$ is closed but 1-degenerate, we shall refer to the pair $(\mathcal{M},\Omega)$ as a pre-multisymplectic manifold. Note that a multisymplectic manifold with $k=2$ is simply a standard symplectic manifold. From a field-theoretic point of view, a particularly relevant class of multisymplectic manifold are the first jet bundles over a given ($n$-dimensional differentiable) base manifold, endowed with the Poincaré-Cartan $(n+1)$-forms [^1] The construction and utility of these manifolds shall become clear in due course. [31].
Given a multisymplectic manifold $(\mathcal{M},\Omega)$, sections of the $k^{\textrm{th}}$ exterior power of the tangent bundle $T\mathcal{M}$ define a class of objects known as multivector fields of degree $k$ (or simply $k$-multivector fields) on $\mathcal{M}$. We denote the space of all such multivector fields $\mathfrak{X}^{k}(\mathcal{M}):=\Gamma(\mathcal{M},\wedge^{k}\,T\mathcal{M})$. A $k$-multivector field $\boldsymbol{X}\in\mathfrak{X}^{k}(\mathcal{M})$ is said to be locally decomposable if, for every point $p\in\mathcal{M}$, there exists an open neighbourhood $\mathcal{U}_{p}\subset\mathcal{M}$, and vector fields $X_{1},\,\cdots\,,X_{k}\in\mathfrak{X}^{\infty}(\mathcal{U}_{p})$, such that
$\boldsymbol{X}|_{\mathcal{U}_{p}}=X_{1}\wedge\,\cdots\,\wedge X_{k}$
We say that an $m$-dimensional distribution $\mathcal{D}\subset T\mathcal{M}$ is locally associated to a non-zero $\boldsymbol{X}\in\mathfrak{X}^{m}(\mathcal{M})$, if there exists some connected open set $\mathcal{V}\subset\mathcal{M}$, such that $\boldsymbol{X}|_{\mathcal{V}}\in\Gamma(\mathcal{V},\wedge^{m}\mathcal{D}|_{\mathcal{V}})$. Note that it is possible for two multivector fields $\boldsymbol{X},\boldsymbol{Y}\in\mathfrak{X}^{m}(\mathcal{M})$ to be locally associated to the same distribution; in such a case, there exists some function $f\in C^{\infty}(\mathcal{V})$, defined on the common connected open subset $\mathcal{V}$, such that $\boldsymbol{Y}=f\boldsymbol{X}$. We shall declare any such multivector fields to be equivalent, and henceforth work with the equivalence classes
$\{\boldsymbol{X}\}_{\mathcal{V}}=\{\boldsymbol{Y}\in\Gamma(\wedge^{m}\mathcal{D}|_{\mathcal{V}})\;|\;\boldsymbol{Y}=f\boldsymbol{X}\;;\;f\in C^{\infty}(\mathcal{V})\,\}$
We refer to a multivector field $\boldsymbol{X}\in\mathfrak{X}^{m}(\mathcal{M})$ as integrable if its locally associated distribution $\mathcal{D}$ is integrable; note that all other multivector fields belonging to the equivalence class of $\boldsymbol{X}$ are automatically integrable, and further, they possess the same integral manifolds [32].
When studying classical field theories, the cases of greatest geometrical interest are those in which we have some bundle structure $\pi:E\rightarrow M$; often this will correspond to the first jet bundle, with $M$ representing the spacetime manifold; however, we shall, for now, retain a somewhat more general setup. We are then interested in the case when the integral manifolds of a given distribution coincide with sections of $\pi$. By this, we mean that, given a section $s:M\rightarrow E$, we shall consider only those integral manifolds $\mathcal{S}\subset E$ which may be expressed as $\mathcal{S}=s(U)$, with $U\subset M$.
Intuitively, this requirement gives rise to a set of integral manifolds which ‘cover’ the total space, in a manner which is transverse to each of the fibres. This notion is made somewhat more precise by introducing the concept of $\pi$-transversality: a multivector field $\boldsymbol{X}\in\mathfrak{X}^{m}(E)$ is said to be $\pi$-transverse if, at each point $p\in E$
$\iota_{\scriptscriptstyle\boldsymbol{X}}(\pi^{*}\xi)=0$
for all $\xi\in\Omega^{m}(M)$ which satisfy $\xi(\pi(p))\neq 0$. Here, the contraction between a locally decomposable $m$-multivector field $\boldsymbol{X}=X_{1}\wedge\,\cdots\,\wedge X_{m}$ and a differential $k$-form $\Xi\in\Omega^{k}(E)$ is naturally defined as
$\iota_{\scriptscriptstyle\boldsymbol{X}}\Xi=\begin{cases}\iota_{\scriptscriptstyle X_{m}}\,\cdots\;\iota_{\scriptscriptstyle X_{1}}\,\Xi\quad&\textrm{if $m\leqslant k$}\\ 0&\textrm{if $m>k$}\end{cases}$
The key statement relating to $\pi$-transverse multivector fields is the following: if $\boldsymbol{X}\in\mathfrak{X}^{m}(E)$ is a locally decomposable integrable multivector field, then it is $\pi$-transverse if and only if its integral manifolds are local sections of $\pi$ [30, 32].
II.2 Lagrangian Field Theory
Having introduced the majority of the mathematical background pertaining to multisymplectic geometry, we shall now illustrate how these ideas are implemented in the Lagrangian description of classical field theories. The fundamental elements of such theories comprise a fibre bundle $\pi:E\rightarrow M$ over the spacetime manifold $M$; we take $M$ to be $m$-dimensional and orientatable, with volume form $\omega$. The total space $E$ is then an $(n+m)$-dimensional manifold, which we refer to as the covariant configuration space; for the vast majority of relevant examples, it will be sufficient to consider total spaces of the form $E=M\times C$, in which $C$ is identified as the $n$-dimensional ‘field space’. Finally, in addition to the bundle $\pi:E\rightarrow M$, we also require the first jet bundle $\kappa:J^{1}E\rightarrow E$ of sections of $\pi$ [33, 34]; this becomes a manifold over $M$ upon introducing the map $\widehat{\pi}:=\pi\circ\kappa:J^{1}E\rightarrow M$. The dimension of $J^{1}E$ is $m+n+nm$, and local coordinates adapted to the bundle structure will be written as $(x^{\mu},y^{a},y^{a}_{\mu})$, with $1\leqslant\mu\leqslant m$ and $1\leqslant a\leqslant n$. Often, when our spacetime manifold is equipped with a Lorentzian metric, we shall adopt the alternative convention that $\textrm{dim}\,M=d$, and $0\leqslant\mu\leqslant d-1$; this will be made clear within the context of the relevant section.
These geometrical elements allow us to define a Lagrangian system on $J^{1}E$, commencing with the introduction of the Lagrangian density $\mathcal{L}$, which is a smooth bundle morphism
$\mathcal{L}:J^{1}E\;\longrightarrow\;\wedge^{m}M$
Since $M$ is orientable, we may consider the Lagrangian density to be a $\widehat{\pi}-$semibasic $m$-form on $J^{1}E$. In general, a differential form $\xi\in\Omega^{m}(J^{1}E)$ is said to be $\widehat{\pi}-$semibasic if $\xi(X_{1},\,\cdots\,,X_{m})=0$ when any one of the vectors $X_{i}\in\mathfrak{X}^{\infty}(J^{1}E)$ belongs to the vertical subspace of $\widehat{\pi}$, that is when $X_{i}\in\textrm{ker}\,\widehat{\pi}_{*}$ [35]. In local bundle coordinates, we shall write
$\mathcal{L}(x^{\mu},y^{a},y^{a}_{\mu})=L(x^{\mu},y^{a},y^{a}_{\mu})\,\widehat{\pi}^{*}\omega$ (1)
in which $L:J^{1}E\rightarrow\mathbb{R}$ is referred to as the Lagrangian function, and $\widehat{\pi}^{*}\omega$ is the volume form on $M$, pulled back to $J^{1}E$.
In complete analogy to the tangent bundle, upon which we study the standard Lagrangian formulation of particles [36, 37], the space $J^{1}E$ is endowed with a canonical structure: the vertical endomorphism [38, 39, 40]; this is used to define the Cartan forms $\Theta_{L}\in\Omega^{m}(J^{1}E)$ and $\Omega_{L}\in\Omega^{m+1}(J^{1}E)$ [41]. In local bundle coordinates $(x^{\mu},y^{a},y^{a}_{\mu})$, we have
$\Theta_{L}=\frac{\partial L}{\partial y^{a}_{\mu}}\,dy^{a}\wedge d^{m-1}x_{\mu}-\left(\frac{\partial L}{\partial y^{a}_{\mu}}\,y^{a}_{\mu}-L\right)\,d^{m}x\quad\quad\quad\quad\Omega_{L}=-\,d\Theta_{L}$ (2)
The quantity $d^{m-1}x_{\mu}$ is defined in the obvious manner as $d^{m-1}x_{\mu}:=\iota_{\scriptscriptstyle\partial_{\mu}}d^{m}x$. We shall then call the pair $(J^{1}E,\Omega_{L})$ a Lagrangian system, and further specify this to be regular if $\Omega_{L}$ is multisymplectic, and singular if $\Omega_{L}$ is pre-multisymplectic [42].
From the Poincaré-Cartan $m$-form (2), we define the objects $\theta_{L}^{\mu}$ according to
$\theta^{\mu}_{L}=-\,\iota_{\partial_{m}}\,\cdots\;\iota_{\partial_{1}}(\Theta_{L}\wedge dx^{\mu})=\frac{\partial L}{\partial y^{\mu}_{a}}\,dy^{a}$ (3)
Note that these are 1-forms which carry a spacetime index; the $m$-form $\Theta_{L}$ may now be written more succinctly as
$\Theta_{L}=\theta_{L}^{\mu}\wedge d^{m-1}x_{\mu}-E_{L}\,d^{m}x$ (4)
in which $E_{L}\in C^{\infty}(J^{1}E)$ is the energy function corresponding to the Lagrangian $\mathcal{L}$.
The field equations for a Lagrangian system $(J^{1}E,\Omega_{L})$ may be derived via a variational principle (details of which can be found in [31], for example); in the current context however, it suffices to acknowledge that this variational principle seeks a critical section $\phi\in\Gamma(M,E)$ determined by the following criterion
$\phi\;\,\textrm{critical}\quad\quad\Longleftrightarrow\quad\quad\frac{d}{dt}\biggr|_{t=0}\,\int_{M}(j^{1}\phi_{t})\Theta_{L}=0$
Here the variations $\phi_{t}=\sigma_{t}\hskip 0.19919pt\circ\hskip 0.19919pt\phi$ are constructed using a one-parameter family of transformations which generate any compactly-supported vector field in the vertical bundle $\mathfrak{X}^{V(\pi)}(E)$. The section $\phi$ may be represented in local coordinates as $\phi(x)=(x^{\mu},y^{a}(x))$. In the above integral, we are interested in the canonical lifting (also referred to as the first jet prolongation) of $\phi$ to $J^{1}E$, denoted $j^{1}\phi$; this is given locally by
$j^{1}\phi(x)=\left(x^{\mu},y^{a}(x),\frac{\partial y^{a}}{\partial x^{\mu}}\biggr|_{x}\,\right)$ (5)
For the purpose of analysing dynamical similarities within a field-theoretic framework, we shall be more interested in a geometrical formulation of the equations of motion, which is entirely equivalent to the above variational approach. In order to present this geometrical formalism, we must expand upon a number of points from section (II.1). Given an $m$-dimensional, $\widehat{\pi}-$transverse, integrable distribution $\mathcal{D}$, we know that it is possible to associate to it an equivalence class of multivector fields $\{\boldsymbol{X}\}$; in local adapted coordinates, a representative of this equivalence class may be expressed as
$\boldsymbol{X}=\bigwedge\limits_{\mu=1}^{m}f\left(\frac{\partial}{\partial x^{\mu}}+F_{\mu}^{a}\,\frac{\partial}{\partial y^{a}}+K_{\mu\nu}^{a}\,\frac{\partial}{\partial y^{a}_{\nu}}\right)$ (6)
in which $f\in C^{\infty}(J^{1}E)$ is some arbitrary (but non-vanishing) function. In almost all practical examples, the condition of $\widehat{\pi}-$transversality is most readily imposed by demanding that $\iota_{\scriptscriptstyle\boldsymbol{X}}(\widehat{\pi}^{*}\omega)=1$, which sets $f=1$ in the above expression.
Suppose that $\psi\in\Gamma(M,J^{1}E)$ is a section of $\widehat{\pi}$, which we choose to represent locally as $\psi(x)=(x^{\mu},y^{a}(x),y^{a}_{\mu}(x))$; we shall refer to this as an integral section of the multivector field $\boldsymbol{X}$ if $F_{\mu}^{a}=y^{a}_{\mu}$ in the expansion (6). Additionally, an integral section is said to be holonomic if it is the canonical lifting of a section $\phi\in\Gamma(M,E)$, that is to say, if we may write $\psi=j^{1}\phi$. A locally decomposable multivector field is then called holonomic if its integral sections are all holonomic. If a locally decomposable multivector field $\boldsymbol{X}$ has the form (6), but is not integrable, then we say that it is semi-holonomic.
The above discussion is quite general; let us now see how these ideas apply in the context of the Lagrangian field equations. Suppose that $\phi\in\Gamma(M,E)$ is a critical section for the Lagrangian system $(J^{1}E,\Omega_{L})$, as determined by the variational principle. Then the canonical lifting $j^{1}\phi$ satisfies
$(j^{1}\phi)^{*}\,\iota_{\scriptscriptstyle X}\Omega_{L}=0\quad\quad\quad\textrm{for all}\;X\in\mathfrak{X}^{\infty}(J^{1}E)$ (7)
Additionally, the section $j^{1}\phi$, expressed in local coordinates as $j^{1}\phi(x)=\left(x^{\mu},y^{a}(x),\frac{\partial y^{a}}{\partial x^{\mu}}|_{x}\right)$, satisfies the following Euler-Lagrange field equations
$\frac{\partial}{\partial x^{\mu}}\,\left(\frac{\partial L}{\partial y^{a}_{\mu}}\circ j^{1}\phi\right)-\frac{\partial L}{\partial y^{a}}\circ j^{1}\phi=0$ (8)
Finally, we find that $j^{1}\phi$ is an integral section of a class of holonomic multivector fields $\{\boldsymbol{X}_{L}\}$, all of which satisfy
$\iota_{\scriptscriptstyle\boldsymbol{X}_{L}}\Omega_{L}=0$ (9)
Note that equations (8) and (9) are highly reminiscent of their standard symplectic analogues [43, 44]. Indeed, this must be the case, given that, as mentioned in section (II.1), symplectic manifolds are but a special case of multisymplectic manifolds, for which the 1-non-degenerate form $\Omega$ coincides with the symplectic 2-form [45].
II.3 Hamiltonian Field Theory
Having discussed the Lagrangian formulation of multisymplectic field theory, we would now like to consider the Hamiltonian counterpart. The construction of a covariant Hamiltonian formalism is less straightforward, since there is no uniquely-defined multiphase space [46, 47, 48]. In what follows, we shall adopt the simplest and most ubiquitous framework found throughout the literature, and construct the Hamiltonian formalism using the so-called extended and restricted multimomentum bundles [49, 50].
Let us denote by $\wedge^{m}_{2}\,T^{*}E$ the space of differential $m$-forms on $E$ that vanish when acting upon two $\pi$-vertical vector fields. We shall refer to $\wedge^{m}_{2}\,T^{*}E$ as the extended multimomentum bundle; its local coordinates adapted to the underlying bundle structure are $(x^{\mu},y^{a},p^{\mu}_{a},p)$, and there exist two natural submersions
$\mu:\wedge^{m}_{2}\,T^{*}E\rightarrow E\quad\quad\textrm{and}\quad\quad\bar{\mu}=\pi\circ\mu:\wedge^{m}_{2}\,T^{*}E\rightarrow M$
The extended multimomentum bundle $\wedge^{m}_{2}\,T^{*}E$ admits a differential $m$-form $\Theta$, known as the Liouville form, which, in local coordinates, is expressed as
$\Theta=p^{\mu}_{a}\,dy^{a}\wedge d^{m-1}x_{\mu}+p\,d^{m}x$ (10)
This Liouville $m$-form then allows us to define a multisymplectic form $\Omega=-\,d\Theta\in\Omega^{m+1}(\wedge^{m}_{2}\,T^{*}E)$; however, the extended multimomentum bundle $\wedge^{m}_{2}\,T^{*}E$ is not the arena in which multisymplectic Hamiltonian dynamics takes place. In order to define this manifold, we first consider the bundle whose sections are $\pi$-semibasic $m$-forms on $E$, which we denote by $\pi^{*}\hskip-2.56073pt\wedge^{m}\hskip-1.42262ptT^{*}M$, and introduce the quotient
$J^{1}E^{*}:=\frac{\wedge^{m}_{2}\,T^{*}E}{\pi^{*}\hskip-2.56073pt\wedge^{m}\hskip-1.42262ptT^{*}M}$
This space is made into a bundle, which we call the restricted multimomentum bundle, upon defining the following projection maps
$\sigma:J^{1}E^{*}\rightarrow E\quad\quad\quad\quad\bar{\sigma}=\pi\circ\sigma:J^{1}E^{*}\rightarrow M$
The restricted multimomentum bundle will serve as our covariant multiphase space for Hamiltonian dynamics. Local coordinates on $J^{1}E^{*}$, adapted to the bundle structure of $E$, are $(x^{\mu},y^{a},p^{\mu}_{a})$. Additionally, we have a naturally-defined map $\widetilde{\mathcal{FL}}:J^{1}E\rightarrow\wedge^{m}_{2}T^{*}E$, known as the extended Legendre transform; this is then restricted to the map $\mathcal{FL}:=\mu\circ\widetilde{\mathcal{FL}}:J^{1}E\rightarrow J^{1}E^{*}$, which acts on local bundle coordinates according to
$\mathcal{FL}^{*}\,x^{\mu}=\;x^{\mu}\quad\quad\quad\quad\mathcal{FL}^{*}\,y^{a}=y^{a}\quad\quad\quad\quad\mathcal{FL}^{*}\,p^{\mu}_{a}=\frac{\partial L}{\partial y^{a}_{\mu}}$ (11)
In complete analogy to the standard symplectic case, we shall call our Lagrangian system $(J^{1}E,\Omega_{L})$ regular if $\mathcal{FL}$ is a local diffeomorphism, hyperregular if this diffeomorphism is global, and singular if $\mathcal{FL}$ is not a diffeomorphism [51, 52, 53]. Additionally, we shall further specify a singular Lagrangian to be almost-regular [54], if the following three conditions are satisfied:
  • $\mathcal{I}=\mathcal{FL}(J^{1}E)$ is a closed submanifold of $J^{1}E^{*}$
  • The map $\mathcal{FL}$ is a submersion onto its image
  • The fibres $\mathcal{FL}^{-1}(p)$ are connected submanifolds of $J^{1}E$ for all points $p\in\mathcal{I}$
The regular and hyperregular cases are treated in an identical manner, with the exception that the regular case requires us to work locally on $J^{1}E$, rather than globally. The almost-regular case, on the other hand, must be addressed separately, through use of a suitable constraint algorithm [42]; further comments will be made on the treatment of singular Lagrangians in due course, as they are particularly relevant to the discussion of gauge theories within the multisymplectic framework [55, 56].
Pulling back the Liouville forms $\Theta$ and $\Omega$ along a Hamiltonian section $h:J^{1}E^{*}\rightarrow\wedge^{m}_{2}\,T^{*}E$, we obtain corresponding $m$ and $(m+1)$-forms on $J^{1}E^{*}$
$\Theta_{h}=h^{*}\Theta\;\in\;\Omega^{m}(J^{1}E^{*})\quad\quad\quad\textrm{and}\quad\quad\quad\Omega_{h}=-\,d\Theta_{h}\;\in\;\Omega^{m+1}(J^{1}E^{*})$ (12)
We refer to these objects as the Hamilton-Cartan $m$ and $(m+1)$-forms, and further define the pair $(J^{1}E^{*},\Omega_{h})$ to be a multisymplectic Hamiltonian system. The Hamiltonian function $H:\mathcal{U}\subset J^{1}E^{*}\rightarrow\mathbb{R}$ is then defined in terms of the section $h$ according to the following local coordinate expression
$h(x^{\mu},y^{a},p^{\mu}_{a})=\left(x^{\mu},y^{a},p^{\mu}_{a},p=-\,H(x^{\nu},y^{b},p^{\nu}_{b})\right)$ (13)
With this definition of the Hamiltonian function, the Hamilton-Cartan forms $\Theta_{h}$ and $\Omega_{h}$ are expressed locally as
$\begin{split}\Theta_{h}&=p^{\mu}_{a}\,dy^{a}\wedge d^{m-1}x_{\mu}-H\,d^{m}x\\ \vskip 0.56905pt\Omega_{h}=-\,&dp^{\mu}_{a}\wedge dy^{a}\wedge d^{m-1}x_{\mu}+dH\wedge d^{m}x\end{split}$ (14)
From $\Omega_{h}$, we shall also define the following 2-forms
$\omega_{h}^{\mu}:=-\,\iota_{\partial_{m}}\cdots\,\iota_{\partial_{1}}(\Omega_{h}\wedge dx^{\mu})=dy^{a}\wedge\,dp^{\mu}_{a}$ (15)
which will be used extensively in section (IV.2), when constructing vector fields corresponding to dynamical similarities.
In a similar spirit to the results of section (II.2), the Hamiltonian field equations, like their Lagrangian counterpart, may also be deduced from a variational principle. This so-called Hamilton-Jacobi principle seeks a critical section $\psi\in\Gamma(M,J^{1}E^{*})$, which, in this context is defined as
$\psi\;\,\textrm{critical}\quad\quad\Longleftrightarrow\quad\quad\frac{d}{dt}\biggr|_{t=0}\,\int_{M}\psi_{t}^{*}\Theta_{h}=0$
In a similar manner to the corresponding Lagrangian principle, the variations $\psi_{t}=f_{t}\circ\psi$ are constructed using a one-parameter family of transformations $\{f_{t}\}$ which generate any compactly-supported $\bar{\sigma}$-vertical vector field on $J^{1}E^{*}$.
The equivalent geometrical formulation of the Hamilton-Jacobi principle seeks an $m$-dimensional, $\bar{\sigma}$-transverse, integrable distribution $\mathcal{D}$, whose integral manifolds are critical sections of this variational problem. As in the Lagrangian case, such distributions are associated with classes of locally decomposable, $\bar{\sigma}$-transverse multivector fields $\{\boldsymbol{X}\}$, which may be expressed locally as
$\boldsymbol{X}=\bigwedge\limits_{\mu=1}^{m}\,f\left(\frac{\partial}{\partial x^{\mu}}+F^{a}_{\mu}\,\frac{\partial}{\partial y^{a}}+G^{\nu}_{a\mu}\,\frac{\partial}{\partial p^{\nu}_{a}}\right)$ (16)
for some non-vanishing function $f\in C^{\infty}(J^{1}E^{*})$. Continuing to draw parallels between the Lagrangian and Hamiltonian cases, if $\psi\in\Gamma(M,J^{1}E^{*})$ is a critical section of $\bar{\sigma}$, then
$\psi^{*}\,\iota_{\scriptscriptstyle X}\Omega_{h}=0\quad\quad\quad\quad\textrm{for all}\;X\in\mathfrak{X}^{\infty}(J^{1}E^{*})$ (17)
Additionally, if we express $\psi$ locally as $\psi(x)=(x^{\mu},\psi^{a}(x),\psi^{\mu}_{a}(x))$, since (17) holds for all vector fields $X\in\mathfrak{X}^{\infty}(J^{1}E^{*})$, using the local expression (14) for the Hamilton-Cartan form $\Omega_{h}$, we obtain the following Hamilton De-Donder Weyl equations
$\frac{\partial\psi^{a}}{\partial x^{\mu}}\biggr|_{x}=\frac{\partial H}{\partial p^{\mu}_{a}}\biggr|_{\psi(x)}\quad\quad\quad\frac{\partial\psi^{\mu}_{a}}{\partial x^{\mu}}\biggr|_{x}=-\,\frac{\partial H}{\partial y^{a}}\biggr|_{\psi(x)}$ (18)
Finally, $\psi$ is an integral section of a class of $\bar{\sigma}$-transverse integrable multivector fields $\{\boldsymbol{X}_{h}\}$, each of which satisfies
$\iota_{\scriptscriptstyle\boldsymbol{X}_{h}}\Omega_{h}=0$ (19)
We conclude this discussion describing the relationship between the critical sections of each of the two variational principles introduced thus far. Suppose that $\phi\in\Gamma(M,E)$ is a critical section for the (hyperregular) Lagrangian system $(J^{1}E,\Omega_{L})$. The object $\psi=\mathcal{FL}\circ j^{1}\phi\in\Gamma(M,J^{1}E^{*})$ then satisfies the Hamilton-Jacobi variational problem. Conversely, given a critical section $\psi\in\Gamma(M,J^{1}E^{*})$, we introduce $\phi=\sigma\circ\psi$, which is a critical section of the Lagrangian variational problem. These relationships are captured in the following commutative diagram [31]
[…]
II.4 An Example: The Free Scalar Field
The tone we have adopted thus far has been highly formal; in order to illustrate the utility of the multisymplectic formalism, and to provide a less abstract presentation of the ideas we have developed, we shall consider a simple, albeit relevant example: that of a free scalar field in a Minkowski background [57].
Spacetime coordinates on $M=\mathbb{R}^{1,3}$ are denoted $x^{\mu}$, with $\mu=0,1,2,3$; since we have a single real scalar field, the covariant configuration space $E$ is simply $E=M\times\mathbb{R}$, and we shall take local coordinates on $J^{1}E$ to be $(x^{\mu},y,y_{\mu})$. In this way, the Lagrangian function $L:J^{1}E\rightarrow\mathbb{R}$ is given by
$L(x^{\mu},y,y_{\mu})=\frac{1}{2}\eta^{\mu\nu}y_{\mu}y_{\nu}-\frac{1}{2}m^{2}y^{2}$ (20)
in which we have adopted the mostly minus convention $\eta_{\mu\nu}=\textrm{diag}(+1,-1,-1,-1)$, more commonly employed throughout the quantum field theory literature [58, 59]. The scalar energy $E_{L}\in C^{\infty}(J^{1}E)$ for this system is simply
$E_{L}=\frac{\partial L}{\partial y_{\mu}}\,y_{\mu}-L=\frac{1}{2}\eta^{\mu\nu}y_{\mu}y_{\nu}+\frac{1}{2}m^{2}y^{2}$ (21)
The Poincaré-Cartan form is of degree 4, and has the following local coordinate expression
$\Theta_{L}=\eta^{\mu\nu}y_{\nu}\,dy\wedge d^{3}x_{\mu}-\left(\frac{1}{2}\eta^{\mu\nu}y_{\mu}y_{\nu}+\frac{1}{2}m^{2}y^{2}\right)d^{4}x$ (22)
The equations of motion for a holonomic section $\phi(x)=(x^{\mu},y(x))$, calculated according to (8), read
$\eta^{\mu\nu}\partial_{\mu}\partial_{\nu}y-m^{2}y=0$
which, as anticipated, is simply the Klein-Gordon equation for a single scalar field in flat spacetime. We also highlight that the derivatives in (8) are to be evaluated along the canonical lifting of $\phi$, which implies that $y_{\mu}\rightarrow\partial_{\mu}y$.
The momenta $p^{\mu}$ canonically conjugate to the field variable $y$ are given by
$p^{\mu}=\frac{\partial L}{\partial y_{\mu}}=\eta^{\mu\nu}y_{\nu}$
Consequently, the Hamiltonian function reads
$H=\frac{1}{2}\eta_{\mu\nu}p^{\mu}p^{\nu}+\frac{1}{2}m^{2}y^{2}$
Consider a locally decomposable, integrable multivector field $\boldsymbol{X}\in\mathfrak{X}^{4}(J^{1}E^{*})$, expressed in local coordinates as
$\boldsymbol{X}=\bigwedge\limits_{\mu=0}^{3}\left(\frac{\partial}{\partial x^{\mu}}+F_{\mu}\frac{\partial}{\partial y}+G^{\nu}_{\mu}\frac{\partial}{\partial p^{\nu}}\right)$
in which, referring to (16), $\bar{\sigma}$-transversality has been enforced setting the multiplicative function $f$ to unity. Using the following expression for the Hamilton-Cartan 5-form
$\Omega_{h}=-\,dp^{\mu}\wedge dy\wedge d^{3}x_{\mu}+(\eta_{\mu\nu}p^{\mu}\,dp^{\nu}+m^{2}y\,dy)\,\wedge\,d^{4}x$
it follows that the coordinate-free equations of motion (19) become
$F_{\mu}=\frac{\partial H}{\partial p^{\mu}}=\eta_{\mu\nu}p^{\nu}\quad\quad\quad\quad G^{\mu}_{\mu}=-\,\frac{\partial H}{\partial y}=-\,m^{2}y$
Supposing that this holonomic multivector field $\boldsymbol{X}$ has integral sections of the form $\psi(x)=\left(x^{\mu},y(x),p^{\mu}(x)\right)$, we may write
$\frac{\partial y}{\partial x^{\mu}}=F_{\mu}=\eta_{\mu\nu}p^{\nu}\quad\quad\quad\quad\frac{\partial p^{\mu}}{\partial x^{\mu}}=G^{\mu}_{\mu}=-\,m^{2}y$
and these equations may easily be shown to be equivalent to the Klein-Gordon equation, as deduced from the Lagrangian formalism.
II.5 Singular Lagrangians
The free scalar field theory of the previous example was a particularly simple case of a hyperregular Lagrangian; such functions admit globally-defined multivector field solutions which are locally decomposable, $\widehat{\pi}-$transverse and have holonomic integral sections. Many of the Lagrangian functions encountered in more physically interesting contexts are not of this nature, and require the implementation of a geometric constraint algorithm, as developed in [42, 60], for example.
As we have seen, the differential form $\Theta_{L}$ associated with a singular Lagrangian is, in general, pre-multisymplectic; as such, it is not guaranteed that $\widehat{\pi}-$transverse, integrable multivector field solutions to the equation $\iota_{\scriptscriptstyle\boldsymbol{X}_{L}}\Omega_{L}=0$ exist. Moreover, if such objects do exist, they are, with almost complete certainty, not defined on the entirety of $J^{1}E$. In the most favourable of cases, solutions exist on some submanifold $S_{f}\subset J^{1}E$; this is the final constraint submanifold, obtained through successive iteration of the geometric constraint algorithm.
In practice, this algorithm is implemented locally, following a number of steps, in which we define spaces of successively more restricted multivector fields. In particular, we begin by introducing the space
$\textrm{ker}^{m}\,\Omega_{L}:=\{\boldsymbol{X}_{L}\in\mathfrak{X}^{m}(J^{1}E)\;|\;\iota_{\scriptscriptstyle\boldsymbol{X}_{L}}\Omega_{L}=0\,\}$
consisting of all $m$-multivector fields which are solutions to the equations of motion, but are not necessarily $\widehat{\pi}-$transverse or integrable. Within this space, we then seek the subset of those multivector fields which are locally decomposable and $\widehat{\pi}-$transverse, denoting the resulting subspace $\textrm{ker}^{m}_{\widehat{\pi}}\,\Omega_{L}\subset\textrm{ker}^{m}\,\Omega_{L}$.
Following this, we further restrict $\textrm{ker}^{m}_{\widehat{\pi}}\,\Omega_{L}$ to the subset of semi-holonomic solutions, which amounts to setting $F_{\mu}^{a}=y^{a}_{\mu}$ in the decomposition (6). In general, this step gives rise to further consistency conditions, and possibly to new constraints; in order to ensure that the dynamics of the system remain confined to the constraint submanifold, it is necessary to impose tangency conditions on each of the constraints. This is achieved by demanding that the Lie derivative of any constraint along each of the vector fields which constitute $\boldsymbol{X}_{L}$ vanish.
Finally, we perform an analysis of the integrability of our semi-holonomic solutions, in order to reach the final subspace $\textrm{ker}^{m}_{\scriptscriptstyle H}\,\Omega_{L}$, consisting of locally decomposable, $\widehat{\pi}-$transverse, holonomic multivector fields. Examples of the local implementation of this geometric algorithm may be found in [61] and [56], for example.
Throughout this analysis, we have assumed the singular Lagrangian to be almost-regular; consequently, we know that $\mathcal{I}=\mathcal{FL}(J^{1}E)$ is a closed submanifold of $J^{1}E^{*}$. In such cases, it is possible to define the Hamilton-Cartan forms $\Theta_{h}^{\scriptscriptstyle\mathcal{I}}$ and $\Omega_{h}^{\scriptscriptstyle\mathcal{I}}$ restricted to $\mathcal{I}$, and formulate the Hamilton-Jacobi principle, as above, on the (generally) pre-multisymplectic manifold $(\mathcal{I},\Omega^{\scriptscriptstyle\mathcal{I}}_{h})$. As in the Lagrangian case, the multivector field solutions to the equations of motion are not guaranteed to exist, and in the case that they do, we must apply the above algorithm in order to deduce the appropriate final constraint submanifold of $\mathcal{I}$.
III Multicontact Formulation of Classical Field Theories
Over the course of the preceding sections, we have dedicated significant effort to understanding how multisymplectic geometry may be used to formulate a rigorous description of classical field theories. Since the formalism we have developed generalises the concepts encountered in standard symplectic geometry, it should be relatively clear that all field theories treated within this framework are conservative in nature.
As we shall see, the identification of a dynamical similarity, and subsequent implementation of a symmetry reduction procedure, will inevitably necessitate the introduction of a framework which accommodates systems of a frictional (or dissipative) nature [24, 62, 63, 21]. The most appropriate arena in which to study such a class of field theories is the multicontact formalism [64, 65, 66], whose construction we shall summarise somewhat more succinctly than our treatment of multisymplectic geometry.
III.1 Lagrangian Formalism
Throughout this section, we shall continue to work with the fibred manifolds $\pi:E\rightarrow M$, and $\kappa:J^{1}E\rightarrow E$, with $\textrm{dim}\,M=m$ and $\textrm{dim}\,E=n+m$. The novelty of the multicontact formalism is that the Lagrangian density is now an $m$-form on the manifold
$\mathcal{S}:=J^{1}E\times_{M}\wedge^{m-1}T^{*}M\cong J^{1}E\times\mathbb{R}^{m}$ (23)
The space $\mathcal{S}$ is of dimension $2m+n+nm$, and is a bundle over both $E$, with projection $\tau:\mathcal{S}\rightarrow E$, and $M$, with $\beta=\pi\circ\tau:\mathcal{S}\rightarrow M$. We take local coordinates on $\mathcal{S}$ to be $(x^{\mu},y^{a},y^{a}_{\mu},s^{\mu})$, in which the $s^{\mu}$ span the factor of $\mathbb{R}^{m}$ in (23), and have the field-theoretic interpretation of an action density. Pulling back the volume form $\omega\in\Omega^{m}(M)$ to $\mathcal{S}$, the Lagrangian density is written locally as
$\mathcal{L}(x^{\mu},y^{a},y^{a}_{\mu},s^{\mu})=L(x^{\mu},y^{a},y^{a}_{\mu},s^{\mu})\,\beta^{*}\omega$ (24)
The Lagrangian function $L$ is referred to as regular if the Hessian matrix
$\left(\frac{\partial^{2}L}{\partial y^{a}_{\mu}\partial y^{b}_{\nu}}\right)$
is regular at all points of $\mathcal{S}$. In close analogy to the multisymplectic case, we have the Lagrangian $m$-form $\Theta_{L}$, which, in local coordinates, is given by
$\Theta_{L}=\left(ds^{\mu}-\frac{\partial L}{\partial y^{a}_{\mu}}dy^{a}\right)\,\wedge\,d^{m-1}x_{\mu}+\left(\frac{\partial L}{\partial y^{a}_{\mu}}\,y^{a}_{\mu}-L\right)\,d^{m}x$ (25)
Note that, in contrast to the case of multisymplectic Lagrangian functions, if $L$ is singular, then the associated $m$-form $\Theta_{L}$ is not strictly required to be pre-multicontact, since the corresponding Reeb distribution is not guaranteed to exist. Such pathological examples will not be of interest to us, and so we shall consider any singular Lagrangian to define a pre-multicontact system.
Given an $r$-form $\xi\in\Omega^{r}(\mathcal{S})$, we introduce the following operator
$\begin{split}\bar{d}\;:\;\Omega^{r}(\mathcal{S})\;&\longrightarrow\;\Omega^{r+1}(\mathcal{S})\\ \xi\;&\longmapsto\;\bar{d}\xi:=d\xi+\sigma_{\Theta_{L}}\wedge\,\xi\end{split}$
where $\sigma_{\Theta_{L}}\in\Omega^{1}(\mathcal{S})$ is known as the dissipation form, and is expressed in local coordinates as
$\sigma_{\Theta_{L}}=-\,\frac{\partial L}{\partial s^{\mu}}\,dx^{\mu}$ (26)
Note that, in general, the operator $\bar{d}$ is not nilpotent, so that there exists no sensible way to formulate cohomology groups, unless $d\sigma_{\Theta_{L}}=0$, in which case, we obtain the so-called Lichnerowicz-Jacobi cohomology [67]. Our motivation for introducing this ‘twisted‘ differential operator is that the Lagrangian $(m+1)$-form associated with $\Theta_{L}$ is not (as in the multisymplectic case) defined simply as $\Omega_{L}=-\,d\Theta_{L}$; instead, we have
$\Omega_{L}=\bar{d}\Theta_{L}=d\Theta_{L}-\frac{\partial L}{\partial s^{\mu}}\,dx^{\mu}\wedge\,\Theta_{L}$ (27)
Note that this $(m+1)$-form is defined without the minus sign used in the multisymplectic case.
Within the context of multisymplectic geometry, we define a conserved quantity to be an (m1)(m-1)-form $\zeta^{\mu}\,d^{m-1}x_{\mu}$, such that, on-shell, we have $\partial_{\mu}\zeta^{\mu}=0$ [55, 68]. Action-dependent field theories, by contrast, contain dissipated quantities; in particular, when the equations of motion are satisfied, instead of a conservation law of the form $\partial_{\mu}\zeta^{\mu}=0$, we find that
$\partial_{\mu}\zeta^{\mu}=\frac{\partial L}{\partial s^{\mu}}\zeta^{\mu}$
which is captured precisely by the dissipation form (26). It is interesting to note that all quantities within an action-dependent field theory follow the same dissipation law, and that this is determined by the Lagrangian (or Hamiltonian) function. Additionally, not all dissipation laws correspond to empirically measurable losses: this fact will be of particular relevance when discussing the physical interpretation of the symmetry-reduced models presented in later sections.
Given a regular multicontact Lagrangian system $(\mathcal{S},\Theta_{L})$, the equations of motion for a holonomic section $\Psi\in\Gamma(M,\mathcal{S})$ are given by
$\Psi^{*}\Theta_{L}=0\quad\quad\quad\textrm{and}\quad\quad\quad\Psi^{*}\iota_{Z}\Omega_{L}=0\quad\textrm{for all}\;Z\in\mathfrak{X}^{\infty}(\mathcal{S})$ (28)
Since $\Psi$ is holonomic, it may be expressed locally as
$\Psi(x)=\left(x^{\mu},y^{a}(x),\frac{\partial y^{a}}{\partial x^{\mu}}\biggr|_{x},s^{\mu}(x)\right)$
with which the coordinate-free equations (28) become
$\begin{split}\frac{\partial}{\partial x^{\mu}}\left(\frac{\partial L}{\partial y^{a}_{\mu}}\circ\Psi\right)=\left(\frac{\partial L}{\partial y^{a}}+\frac{\partial L}{\partial y^{a}_{\mu}}\frac{\partial L}{\partial s^{\mu}}\right)\circ\Psi\\ \frac{\partial s^{\mu}}{\partial x^{\mu}}=L\circ\Psi\end{split}$ (29)
We shall refer to these expressions as the Herglotz-Lagrange field equations [69]. As in the multisymplectic case, the dynamical evolution of our system may also be expressed in terms of multivector fields; in the current context, the equations of motion for locally decomposable $\beta$-transverse multivector fields $\boldsymbol{X}_{L}\in\mathfrak{X}^{m}(\mathcal{S})$ are
$\iota_{\scriptscriptstyle\boldsymbol{X}_{L}}\Theta_{L}=0\quad\quad\quad\quad\quad\iota_{\scriptscriptstyle\boldsymbol{X}_{L}}\Omega_{L}=0$ (30)
III.2 Hamiltonian Formalism
Recall that the Hamiltonian formalism of multisymplectic dynamics was developed on the restricted multimomentum bundle $J^{1}E^{*}$; in order to provide an analogous description for dissipative field theories, we shall work on the space
$\mathcal{S}^{*}:=J^{1}E^{*}\times_{M}\wedge^{m-1}T^{*}M\cong J^{1}E^{*}\times\mathbb{R}^{m}$
Introducing the projection $\bar{\tau}:\mathcal{S}^{*}\rightarrow M$, we obtain a bundle over $M$, of dimension $2m+n+nm$, upon which local coordinates are taken to be $(x^{\mu},y^{a},p^{\mu}_{a},s^{\mu})$.
Given a (hyper)regular Lagrangian system $(\mathcal{S},\Theta_{L})$, there exists a map $\mathcal{FL}:\mathcal{S}\rightarrow\mathcal{S}^{*}$, which acts on local coordinates according to
$\mathcal{FL}^{*}\,x^{\mu}=\;x^{\mu}\quad\quad\quad\quad\mathcal{FL}^{*}\,y^{a}=y^{a}\quad\quad\quad\quad\mathcal{FL}^{*}\,p^{\mu}_{a}=\frac{\partial L}{\partial y^{\mu}_{a}}\quad\quad\quad\quad\mathcal{FL}^{*}\,s^{\mu}=s^{\mu}$ (31)
This Legendre map then allows us to define the Hamiltonian $m$-form on $\mathcal{S}^{*}$ as $\Theta_{L}=\mathcal{FL}^{*}\Theta_{H}$; in local coordinates, $\Theta_{H}$ is given by
$\Theta_{H}=\left(ds^{\mu}-p^{\mu}_{a}\,dy^{a}\right)\,\wedge\,d^{m-1}x_{\mu}+H\,d^{m}x$ (32)
in which we have introduced the Hamiltonian function $H:\mathcal{S}^{*}\rightarrow\mathbb{R}$, related to the Lagrangian energy function according to $E_{L}=\mathcal{FL}^{*}H$. When studying the reduction of multisymplectic field theories to their frictional multicontact counterparts, it will be of benefit to work with the following 1-forms
$\eta^{\mu}:=-\,\iota_{\partial_{m}}\cdots\;\iota_{\partial_{1}}\,(\Theta_{H}\wedge dx^{\mu})=ds^{\mu}-p^{\mu}_{a}\,dy^{a}$ (33)
The dissipation form for action-dependent field theories within the Hamiltonian formalism is given by
$\sigma_{\Theta_{H}}=\frac{\partial H}{\partial s^{\mu}}\,dx^{\mu}$ (34)
which is then used to construct the $(m+1)$-form $\Omega_{H}=\bar{d}\Theta_{H}$, as in (27).
A geometric formulation of the field equations may be given in terms of holonomic sections $\psi:M\rightarrow\mathcal{S}^{*}$, or locally decomposable, $\bar{\tau}$-transverse multivector fields $\boldsymbol{X}_{H}\in\mathfrak{X}^{m}(\mathcal{S}^{*})$. In the former case, we have
$\psi^{*}\Theta_{H}=0\quad\quad\quad\textrm{and}\quad\quad\quad\psi^{*}\iota_{Z}\,\Omega_{H}=0\quad\textrm{for all}\;Z\in\mathfrak{X}^{\infty}(\mathcal{S}^{*})$ (35)
whilst the corresponding equations for multivector fields are
$\iota_{\scriptscriptstyle\boldsymbol{X}_{H}}\Theta_{H}=0\quad\quad\quad\textrm{and}\quad\quad\quad\iota_{\scriptscriptstyle\boldsymbol{X}_{H}}\Omega_{H}=0$ (36)
If we decompose the multivector field $\boldsymbol{X}_{H}$ locally as
$\boldsymbol{X}_{H}=\bigwedge\limits_{\mu=1}^{m}\,\left(\frac{\partial}{\partial x^{\mu}}+A^{a}_{\mu}\,\frac{\partial}{\partial y^{a}}+B^{\nu}_{a\mu}\,\frac{\partial}{\partial p^{\nu}_{a}}+C^{\nu}_{\mu}\frac{\partial}{\partial s^{\nu}}\right)$ (37)
then the coordinate-free expressions (36) become
$A_{\mu}^{a}=\frac{\partial H}{\partial p^{\mu}_{a}}\quad\quad\quad B^{\mu}_{a\mu}=-\,\left(\frac{\partial H}{\partial y^{a}}+p^{\mu}_{a}\frac{\partial H}{\partial s^{\mu}}\right)\quad\quad\quad C^{\mu}_{\mu}=p^{\mu}_{a}\frac{\partial H}{\partial p^{\mu}_{a}}-H$ (38)
Let us suppose that $\psi:M\rightarrow\mathcal{S}^{*}$ is an integral section of $\boldsymbol{X}_{H}$; writing $\psi(x)=\left(x^{\mu},y^{a}(x),p^{\mu}_{a}(x),s^{\mu}(x)\right)$, the equations (35) are given locally by
$\frac{\partial y^{a}}{\partial x^{\mu}}\biggr|_{x}=\frac{\partial H}{\partial p^{\mu}_{a}}\biggr|_{\psi(x)}\quad\quad\quad\frac{\partial p^{\mu}_{a}}{\partial x^{\mu}}\biggr|_{x}=-\,\left(\frac{\partial H}{\partial y^{a}}+p^{\mu}_{a}\frac{\partial H}{\partial s^{\mu}}\right)\biggr|_{\psi(x)}\quad\quad\quad\frac{\partial s^{\mu}}{\partial x^{\mu}}\biggr|_{x}=\left(p^{\mu}_{a}\frac{\partial H}{\partial p^{\mu}_{a}}-H\right)\biggr|_{\psi(x)}$ (39)
We shall refer to (38) and (39) as the Herglotz De-Donder Weyl equations for multivector fields and holonomic sections respectively; they are the appropriate modification of the standard Hamilton De-Donder Weyl equations, adapted to describe field theories which are frictional in nature [70]. Note that the last of the expressions in (39) is simply the condition $\partial_{\mu}s^{\mu}=L^{H}$, expressed in canonically conjugate variables.
III.3 The Scalar Field Revisited
In order to conclude our study of multisymplectic and multicontact geometry, we shall revisit the example of section (II.4), in which we considered a single free scalar field in a flat background spacetime. In an attempt to present an example which highlights the features of the preceding sections, but does not introduce further unnecessary complications, we shall content ourselves with a Lagrangian function which differs from that of (20) only by a simple linear term
$L(x^{\mu},y,y_{\mu},s^{\mu})=\frac{1}{2}\eta^{\mu\nu}y_{\mu}y_{\nu}-\frac{1}{2}m^{2}y^{2}+\lambda_{\mu}s^{\mu}$ (40)
in which $\lambda$ is some constant spacetime vector.
The Lagrangian 4-form $\Theta_{L}$ contains an additional contribution from the action density, and so differs somewhat from (22), that we calculated for the free field on $J^{1}E$
$\Theta_{L}=\left(ds^{\mu}-\eta^{\mu\nu}y_{\nu}\,dy\right)\,\wedge\,d^{3}x_{\mu}\,+\,\biggr(\,\underbrace{\frac{1}{2}\eta^{\mu\nu}y_{\mu}y_{\nu}+\frac{1}{2}m^{2}y^{2}-\lambda_{\mu}s^{\mu}}_{=\;E_{L}}\,\biggr)\,d^{4}x$ (41)
The Herglotz-Lagrange equations of motion for a holonomic section $\Psi:M\rightarrow\mathcal{S}=J^{1}E\times\mathbb{R}^{4}$ are given by
$\eta^{\mu\nu}\partial_{\mu}\partial_{\nu}y-m^{2}y=\eta^{\mu\nu}(\partial_{\mu}y)\lambda_{\nu}$
together with the requirement that $\partial_{\mu}s^{\mu}=L$; due to the presence of the additional term on the right-hand side of the above equation, the Lagrangian (40) describes a scalar field subject to a constant damping, whose strength is determined by $\lambda_{\mu}$.
It is a simple exercise to show that the Hamiltonian function $H:\mathcal{S}^{*}=J^{1}E^{*}\times\mathbb{R}^{4}\rightarrow\mathbb{R}$ is given by
$H(x^{\mu},y,p^{\mu},s^{\mu})=\frac{1}{2}\eta_{\mu\nu}p^{\mu}p^{\nu}+\frac{1}{2}m^{2}y^{2}-\lambda_{\mu}s^{\mu}$ (42)
In order to calculate $\Omega_{H}\in\Omega^{5}(\mathcal{S}^{*})$, we require an expression for the dissipation form $\sigma_{\Theta_{H}}$, which, from (34), we deduce to be
$\sigma_{\Theta_{H}}=\frac{\partial H}{\partial s^{\mu}}\,dx^{\mu}=-\lambda_{\mu}\,dx^{\mu}$ (43)
It thus follows that
$\Omega_{H}=d\Theta_{H}+\sigma_{\Theta_{H}}\wedge\,\Theta_{H}=dH\wedge d^{4}x-dp^{\mu}\wedge dy\wedge d^{3}x_{\mu}-\lambda_{\nu}dx^{\nu}\wedge(ds^{\mu}-p^{\mu}dy)\wedge d^{3}x_{\mu}$
Finally, suppose that
$\boldsymbol{X}_{H}=\bigwedge\limits_{\mu=0}^{3}\,\left(\frac{\partial}{\partial x^{\mu}}+A_{\mu}\,\frac{\partial}{\partial y}+B^{\nu}_{\mu}\,\frac{\partial}{\partial p^{\nu}}+C^{\nu}_{\mu}\frac{\partial}{\partial s^{\nu}}\right)$
is a locally decomposable $\bar{\tau}$-transverse multivector field satisfying $\iota_{\scriptscriptstyle\boldsymbol{X}_{H}}\Theta_{H}=0$ and $\iota_{\scriptscriptstyle\boldsymbol{X}_{H}}\Omega_{H}=0$; in this case, the component functions $A_{\mu}$, $B^{\nu}_{\mu}$, and $C^{\nu}_{\mu}$ satisfy
$A_{\mu}=\frac{\partial H}{\partial p^{\mu}}=\eta_{\mu\nu}p^{\nu}\quad\quad B^{\mu}_{\mu}=-\,\left(\frac{\partial H}{\partial y}+p^{\mu}\frac{\partial H}{\partial s^{\mu}}\right)=-m^{2}y+\lambda_{\mu}p^{\mu}\quad\quad C^{\mu}_{\mu}=p^{\mu}\frac{\partial H}{\partial p^{\mu}}-H=\frac{1}{2}\eta_{\mu\nu}p^{\mu}p^{\nu}-\frac{1}{2}m^{2}y^{2}+\lambda_{\mu}s^{\mu}$
IV An Introduction to Dynamical Similarity
In this section, we shall present a brief introduction to the concept of and motivation behind dynamical similarity; in particular, we begin by illustrating how this specific class of symmetry arises within the familiar setting of symplectic geometry, and how the presence of a dynamical similarity points to a redundancy within our mathematical framework, which may ultimately be eliminated. The field-theoretic generalisation of this process of ‘contact reduction’ is the main objective of this work, and is presented in sections (IV.1) and (IV.2).
Consider a general symplectic manifold $(N,\omega)$, with $\textrm{dim}\,N=2n$; suppose also that $X_{H}\in\mathfrak{X}^{\infty}(N)$ is the (unique) Hamiltonian vector field associated with the function $H:N\rightarrow\mathbb{R}$ via
$\iota_{X_{H}}\omega=dH$ (44)
A vector field $D\in\mathfrak{X}^{\infty}(N)$ is said to constitute a scaling symmetry of the Hamiltonian system $(N,\omega,H)$ if [^2] Here, we use $\mathfrak{L}$ to denote the Lie derivative, so as to reserve $\mathcal{L}$ and $L$ for the Lagrangian density and function respectively.
$\mathfrak{L}_{D}\hskip 1.13809pt\omega=\omega\quad\quad\quad\quad\quad\mathfrak{L}_{D}H=\Lambda H$
for some $\Lambda\in\mathbb{R}$, which we refer to as the degree of the scaling symmetry. The first of these properties indicates that scaling symmetries belong to the set of non-strictly canonical transformations of $(N,\omega)$ [71]. Consider now the Lie bracket of $D$ and $X_{H}$, in particular
$\iota_{\scriptscriptstyle[D,X_{H}]}\hskip 0.85358pt\omega=[\mathfrak{L}_{D},\iota_{X_{H}}]\hskip 1.13809pt\omega=(\Lambda-1)\,dH\quad\quad\quad\implies\quad\quad\quad[D,X_{H}]=(\Lambda-1)X_{H}$
Note that the vector field $D$ is not a symmetry of $X_{H}$ in the conventional sense i.e satisfying $D,X_{H}]=0$; instead, it acts to rescale the Hamiltonian vector field by an overall factor, which justifies our use of the term ‘scaling symmetry’.
A non-trivial corollary of this is that there exists an algebra of invariants of the scaling symmetry, whose dynamical evolution is both autonomous and insensitive to its action [24, 21]. In this sense, we may simplistically speak of a scaling symmetry as ‘mapping a physical solution into a second, indistinguishable one’. Having identified a scaling symmetry $D$, it is then possible to construct a space formed from the quotient of $N$ by the (one-dimensional) orbits of $D$; if we declare two points on the same $D$-orbit to be equivalent [^3] Note that this identification is the mathematical realisation of our qualitative statement regarding mapping equivalent physical solutions into one another., and denote the resulting quotient space $N/\sim$, then this space is of dimension 2n12n-1. Further, given a set of assumptions on the action of $D$, which are not overly restrictive (i.e that the one-parameter family of transformations generating $D$ acts freely and properly on $N$) $N/\sim$ is in fact a smooth manifold, with submersion $\pi:N\rightarrow N/\sim$, that admits a contact distribution $\xi:=\pi_{*}\,\textrm{ker}(\iota_{D}\omega)$ [72].
Within the Hamiltonian framework of symplectic mechanics, we typically work on the cotangent bundle $T^{*}Q$ over the $n$-dimensional configuration space $Q$ [73]; we then identify a scaling function $x:T^{*}Q\rightarrow\mathbb{R}$, which is simply a phase space function satisfying $Dx=x$. For many simple examples, $x$ can be chosen to be a single configuration space variable, with the result that the reduced space $T^{*}Q/\sim$ is isomorphic to $T^{*}\widetilde{Q}\times\mathbb{R}$, in which $\widetilde{Q}$ is a configuration space consisting of n1n-1 variables, with $x$ having been excised. Note that although $\textrm{dim}\,T^{*}\widetilde{Q}=2n-2$, we have eliminated only one degree of freedom, as the factor of $\mathbb{R}$ in $T^{*}\widetilde{Q}\times\mathbb{R}$ is spanned by the action $S$, which has now assumed the role of a dynamical variable.
Dynamical similarities are most prevalently encountered in theories in which our mathematical description makes reference (often in the name of calculational simplicity) to a variable associated with the global scale of the system [21]. The dynamical evolution of the set of observables is not affected by this scale parameter, and so may freely be removed from our ontology [74]. The factorisation of the reduced phase space [^4] This is, of course, not the most general final topology of the reduced space; however, it is both illustrative, and sufficiently common so as to warrant particular attention. into $T^{*}\widetilde{Q}\times\mathbb{R}$ gives a clear and intuitive interpretation, which we shall carry forward, when generalising these ideas to a field-theoretic discussion. A theory possessing a scaling symmetry is dynamically-equivalent to a second, simpler theory, consisting of a scale-invariant piece, together with an action-dependent component, which acts in a way that closely resembles friction [62].
The above statement contains two subtleties, which are of particular note. Firstly, the qualifying phrase ‘dynamically-equivalent’ is vital: the reduced theory faithfully reproduces the on-shell, observable dynamics of the original symplectic system; elsewhere, these two theories need not (and most likely will not) coincide. Additionally, we have described the action-dependent theory as exhibiting ‘friction-like’ effects. Typically, when describing a non-conservative system, one imagines that the mechanical energy lost is dissipated into other, empirically-inaccessible degrees of freedom [21, 75]. For a contact-reduced system, by contrast, the changes in energy may either be losses or gains; however, crucially, they arise as a result of the inaccessibility of the global scale variable to an intrinsic observer. For this reason, we employ the phrase ‘friction-like’, so as to ensure a clear distinction between our notion of non-conservative systems, and that of genuinely dissipative phenomena.
IV.1 Contact Reduction of Lagrangian Systems
Having presented a heuristic outline of the concept of contact reduction for theories of particles, we proceed with a detailed description of the analogous procedure within a field-theoretic framework, starting with the Lagrangian formalism. We will then discuss the geometry and topology of the reduced space, and show, through a series of simple examples, how the reduced theory has a particularly elegant physical interpretation, with potential implications for the quantisation procedure.
Let us consider a (regular) Lagrangian system $(J^{1}E,\Theta_{L})$ over the $d$-dimensional spacetime manifold $M$. Suppose, without loss of generality, that the Lagrangian function $L:J^{1}E\rightarrow\mathbb{R}$ depends upon the two field variables $\phi$ and $\psi$, and that, in order to preserve Poincaré invariance, the only dependence upon the spacetime coordinates $x^{\mu}$ is through that of $\phi(x)$ and $\psi(x)$.
We now identify a transformation $(x^{\mu},\phi,\psi)\mapsto(\lambda^{A}x^{\mu},\lambda^{B}\phi,\psi)$ on the covariant configuration space, under which the Lagrangian rescales as $L\mapsto\lambda^{\Lambda}L$ for $\Lambda\in\mathbb{R}$. This transformation may be implemented via the vector field
$\bar{\Sigma}=Ax^{\mu}\frac{\partial}{\partial x^{\mu}}+B\phi\frac{\partial}{\partial\phi}$ (45)
However, note that $\bar{\Sigma}\in\mathfrak{X}^{\infty}(E)$ is a vector field over $E$, and is therefore not directly applicable to the Lagrangian function. Consequently, we look to reproduce the effects of $\bar{\Sigma}$ by lifting it to a vector field $\Sigma$ on the first jet bundle [76]. It is relatively easy to see that
$\Sigma=B\phi\frac{\partial}{\partial\phi}+(B-A)\phi_{\mu}\frac{\partial}{\partial\phi_{\mu}}-A\psi_{\mu}\frac{\partial}{\partial\psi_{\mu}}$ (46)
correctly reproduces the transformation at the level of the jet bundle. Note that we have assumed that a single field $\phi$ undergoes a rescaling, whilst $\psi$ remains invariant. This may seem somewhat restrictive; however, it is always possible to perform a suitable change of coordinates, such that only one of the fields belonging to $E$ is rescaled; indeed, we shall see an example of this below.
Having deduced the vector field (46), we must then verify that $\Sigma$ satisfies
$\mathfrak{L}_{\Sigma}\,\theta_{L}^{\mu}=\theta^{\mu}_{L}\quad\quad\quad\quad\textrm{for all $\mu=0,\,\cdots\,,d-1$}$ (47)
where, we recall, the $\theta_{L}^{\mu}$ are obtained from the Poincaré-Cartan form $\Theta_{L}$ according to (3). From our discussion above, it should be relatively clear that if $\Sigma$ is a scaling symmetry, then so too is any non-zero multiple of $\Sigma$. Consequently, the notion of a scaling symmetry is defined only up to an overall normalisation, and our choice (47) fixes this ambiguity.
In order to progress further with the reduction process, consider a Herglotz Lagrangian $L^{H}$, embedded within a multisymplectic manifold, of one dimension higher. We shall suppose that the Herglotz Lagrangian depends upon the $k$ scalar fields $\phi^{a}$ ($a=1,\,\cdots,k$), and possibly their first derivatives $\phi^{a}_{\mu}$, together with the action density $s^{\mu}$. Finally, we introduce a new field variable $\rho$, defined in such a way that
$\rho_{\mu}=-\,\frac{\partial L^{H}}{\partial s^{\mu}}$ (48)
Writing our multisymplectic Lagrangian $L$ in the following form
$L=e^{\rho}(L^{H}+\rho_{\mu}s^{\mu})$ (49)
a short calculation reveals that, when restricted to the reduced space, the equations of motion of $L^{H}$ and those of $L$ describe the same dynamics for all observable degrees of freedom.
In light of this, suppose that we have identified a scaling symmetry $\Sigma$ of degree $\Lambda$, as outlined above. Referring to the decomposition (46), we then introduce the scaling function $\xi=\phi^{1/B}$, and, with the coordinate transformation $dy^{\mu}=\xi^{-A}\,dx^{\mu}$, the vector field $\Sigma$ is rendered of the form
$\Sigma=\xi\frac{\partial}{\partial\xi}+\xi_{\bar{\mu}}\frac{\partial}{\partial\xi_{\bar{\mu}}}$ (50)
in which $\xi_{\bar{\mu}}$ denotes the velocity coordinate of $\xi$, taken with respect to $y$, that is $\xi_{\bar{\mu}}=\partial\xi/\partial y^{\mu}$. Finally, making the identification $\xi=e^{\rho}$, we find that the Lagrangian adopts the following form [63]
$L(\rho,\rho_{\bar{\mu}},\psi,\psi_{\bar{\mu}})=e^{\Lambda\rho}f(\rho_{\bar{\mu}},\psi,\psi_{\bar{\mu}})$ (51)
for some function $f$ which, crucially, does not depend upon $\rho$. In these coordinates, the scaling symmetry is simply $\Sigma=\partial_{\rho}$.
In section (II.2), we presented a variational principle used to derive the Euler-Lagrange field equations; this method utilised the Poincaré-Cartan form $\Theta_{L}$, which is somewhat removed from the conventional construction of a variational principle, in which we carry out an integral of the Lagrangian. The origin of this apparent disparity lies in a formal feature of the multisymplectic variational principle, which is rarely discussed within the standard symplectic framework. The variational principle for multisymplectic systems makes use of objects known as Lepagean equivalents [77, 78]. In particular, the equations of motion are, in principle, deduced from the extremisation of the action
$\int_{M}\mathcal{L}\circ j^{1}\phi$
in which $\mathcal{L}\in\Omega^{d}(J^{1}E)$ is the Lagrangian density introduced in (1). However, this is a constrained calculation, since we are obliged to consider only those variations which lie within the space of holonomic sections [49]. It is convenient, therefore, to construct an equivalent variational principle, which allows the full space of sections to be considered without constraints. This is achieved through use of the form $\Theta_{L}$, which is a Lepagean equivalent of the Lagrangian density [79, 80].
Our interest in this seemingly inconsequential mathematical formality lies in the fact that, even though we use the Lepagean equivalent formulation for practical calculations, the equations of motion are, fundamentally, derived from an integral of the Lagrangian density. Consequently, should we wish map physical solutions into other, indistinguishable solutions, we must consider $\mathcal{L}$, and not just the Lagrangian function. This is of particular importance, given that we have made a reparameterisation of the spacetime coordinates, which has a direct effect on the volume form $\omega$. In particular, for a $d$-dimensional spacetime $M$, the redefinition $dy^{\mu}=\xi^{-A}\,dx^{\mu}$ leads to a corresponding change of $\omega\mapsto\xi^{A\cdot d}\,\omega$. We find that the relationship between the coordinate scaling $A$, the degree of the symmetry $\Lambda$, and $d$ is $\Lambda+A\cdot d=1$; this is a generalisation of the result of the time-reparameterisation of symplectic systems [72, 63], and the factor of the spacetime dimension enters the expression due to our reparameterising each of the $d$ coordinates equally.
The conclusion to be drawn from this discussion is that, in order to faithfully map between indistinguishable physical solutions, we must work with the $d$-form $L\,d^{d}x$, and not simply the Lagrangian function in isolation. From (51), we see that
$L\,d^{d}x\;\longrightarrow\;e^{\rho}f(\rho_{\bar{\mu}},\psi,\psi_{\bar{\mu}})\,d^{d}y:=\widehat{L}\,d^{d}y$ (52)
The Lagrangian function $\widehat{L}$ is then of a form reminiscent of (49); indeed, let us consider the Euler-Lagrange field equations (8) for $\rho$, applied to a Lagrangian of the form $\widehat{L}=e^{\rho}f(\rho_{\bar{\mu}},\psi,\psi_{\bar{\mu}})$; for ease of notation, we shall revert to $x^{\mu}$ to label spacetime coordinates, and refrain from placing bars over derivative indices.
$0=\frac{\partial}{\partial x^{\mu}}\left[e^{\rho}\frac{\partial f}{\partial\rho_{\mu}}\right]-e^{\rho}f=e^{\rho}\left[\rho_{\mu}\frac{\partial f}{\partial\rho_{\mu}}+\frac{\partial}{\partial x^{\mu}}\frac{\partial f}{\partial\rho_{\mu}}-f\right]$
$\implies\quad\quad f=\rho_{\mu}\frac{\partial f}{\partial\rho_{\mu}}+\frac{\partial}{\partial x^{\mu}}\frac{\partial f}{\partial\rho_{\mu}}$
Comparing this expression for $f$ to the Lagrangian $L=e^{\rho}(L^{H}+\rho_{\mu}s^{\mu})$, and recalling that we require $\partial_{\mu}s^{\mu}=L^{H}$, it is clear that we should identify the action density and Herglotz Lagrangian as follows [^5] Strictly speaking, $\widehat{L}$ and $L^{H}$ belong to different spaces - the former being a function on $J^{1}E$, and the latter a function on $J^{1}E^{\prime}\times\mathbb{R}^{d}$ - and this should be reflected through appropriate use of submersions between the total space $E$ and the symmetry-reduced space $E^{\prime}$; however, for the sake of clarity, we have chosen to omit such maps.
$s^{\mu}=\frac{\partial f}{\partial\rho_{\mu}}\quad\quad\quad\quad L^{H}=f-\rho_{\mu}s^{\mu}$ (53)
In this construction, we have examined only the equation of motion for $\rho$; however, if we wish to affirm that our symmetry-reduced theory is dynamically-equivalent to that described by the original multisymplectic Lagrangian $\widehat{L}$, we must also verify that the equations of motion of the unscaled variables are faithfully reproduced. In order to do so, we should take into account the fact that, within the symmetry-reduced space, the velocity coordinate $\rho_{\mu}$ is to be replaced with the appropriate expression, in terms of $\psi$, $\psi_{\mu}$, and $s^{\mu}$ [25]. Thus, in the Herglotz-Lagrange equation for $\psi$
$\frac{\partial}{\partial x^{\mu}}\frac{\partial L^{H}}{\partial\psi_{\mu}}-\frac{\partial L^{H}}{\partial\psi}-\frac{\partial L^{H}}{\partial\psi_{\mu}}\frac{\partial L^{H}}{\partial s^{\mu}}=0$ (54)
we have the following terms
$\begin{split}\frac{\partial L^{H}}{\partial\psi_{\mu}}=\frac{\partial f}{\partial\psi_{\mu}}+\frac{\partial f}{\partial\rho_{\nu}}\frac{\partial\rho_{\nu}}{\partial\psi_{\mu}}-s^{\nu}\frac{\partial\rho_{\nu}}{\partial\psi_{\mu}}\\ \frac{\partial L^{H}}{\partial\psi}=\frac{\partial f}{\partial\psi}+\frac{\partial f}{\partial\rho_{\nu}}\frac{\partial\rho_{\nu}}{\partial\psi}-s^{\nu}\frac{\partial\rho_{\nu}}{\partial\psi}\\ \frac{\partial L^{H}}{\partial s^{\mu}}=\frac{\partial f}{\partial\rho_{\nu}}\frac{\partial\rho_{\nu}}{\partial s^{\mu}}-\rho_{\mu}-s^{\nu}\frac{\partial\rho_{\nu}}{\partial s^{\mu}}\end{split}$
Combining these results, and recalling from (53) that $s^{\mu}=\partial f/\partial\rho_{\mu}$, we find that the left-hand side of (54) becomes
$\frac{\partial}{\partial x^{\mu}}\frac{\partial f}{\partial\psi_{\mu}}-\frac{\partial f}{\partial\psi}+\rho_{\mu}\frac{\partial f}{\partial\psi_{\mu}}$
Upon multiplying this expression by $e^{\rho}$, it is apparent that the Herglotz-Lagrange equation (54) implies that
$\frac{\partial}{\partial x^{\mu}}\left(e^{\rho}\frac{\partial f}{\partial\psi_{\mu}}\right)-e^{\rho}\frac{\partial f}{\partial\psi}=0$
which is precisely the multisymplectic field equation for $\psi$. Consequently, we have demonstrated that the Herglotz Lagrangian calculated according to (53) correctly reproduces the dynamics of the original system; moreover, this new description makes no reference to $\rho$, confirming our statement that this scaling variable plays no role in the dynamical evolution of the physically-observable degrees of freedom. The symmetry-reduced system is then multicontact or pre-multicontact, depending upon the Lagrangian form $\Theta_{L^{H}}$; a detailed discussion of the geometry of the symmetry-reduced space is postponed until section (IV.2), having presented both the Lagrangian and Hamiltonian formalisms.
We conclude this section with a number of remarks: firstly, the procedure we have outlined above is often far more straightforward in practice. The vast majority of field theories of interest, which admit dynamical similarities, such as scalar field theories, bosonic string theory, and even General Relativity, do not require spacetime coordinate reparameterisations; in this case, we modify the reduction process slightly. The vector fields (45) and 46) remain valid, setting $A=0$; we also seek a scaling function $\xi=\phi^{1/B}$, as before; however, instead of identifying $\xi=e^{\rho}$, we must take $\xi=e^{\rho/\Lambda}$. This serves to ensure that the Lagrangian acquires the final form $L=e^{\rho}f(\rho_{\mu},\psi,\psi_{\mu})$, which is essential for the subsequently-derived Herglotz Lagrangian to produce the correct on-shell dynamics.
Finally, we must also address the case of degenerate Lagrangians [42, 81]; one of our principal objectives in developing this mathematical framework is the study of classical General Relativity. It has been demonstrated that, upon making a conformal decomposition of the spacetime metric, the Einstein-Hilbert Lagrangian admits a scaling symmetry, in which the scaling variable is precisely the conformal factor [25]. This analysis was performed within the second-order formalism, framed in the language of $k$-symplectic geometry [82]. It will be of great interest, therefore, to examine the details of this dynamical similarity in the first-order vielbein formalism [83, 84]. In addition to General Relativity, any other gauge theory we may wish to study will also require the development of suitable geometric tools, adapted to the treatment of singular Lagrangians [85]. The symmetry reduction of such theories is an area of active research; it is, at present, not clear, for example, if the process of gauge-fixing commutes with the elimination of the scaling variable. Additionally, there remain several open questions regarding the topology of the reduced space, when the original pre-multisymplectic theory is confined to a constraint submanifold of the covariant phase space.
IV.2 Contact Reduction of Hamiltonian Systems
In the previous section, we presented an algorithmic procedure for the symmetry reduction of Lagrangian systems admitting a dynamical similarity; we shall now provide the complementary Hamiltonian construction, together with a detailed discussion of the geometry of the reduced space, within both the Lagrangian and Hamiltonian formalisms. The reduction procedure for the standard symplectic case has been formulated rigorously in [72], and our objective is to generalise these principles to a field-theoretic context.
Suppose that $(J^{1}E^{*},\Omega_{h})$ is a multisymplectic Hamiltonian system (over the $d$-dimensional spacetime manifold $M$), with corresponding Hamiltonian function $H:J^{1}E^{*}\rightarrow\mathbb{R}$. For simplicity, we shall also temporarily assume that the Legendre map is a global diffeomorphism, so that our system is hyperregular. Recall, from section (II.3), that we constructed the 2-forms [^6] Here, we have adapted our notation, so as to reflect the fact that spacetime indices assume values $\mu=0,\,\cdots,d-1$.
$\omega^{\mu}_{h}=-\,\iota_{\partial_{d-1}}\cdots\;\iota_{\partial_{0}}(\Omega_{h}\wedge dx^{\mu})$ (55)
We shall declare a vector field $D\in\mathfrak{X}^{\infty}(J^{1}E^{*})$ to constitute a dynamical similarity of the Hamiltonian system if
$\mathfrak{L}_{D}H=\Gamma H\quad\quad\quad\textrm{and}\quad\quad\quad\mathfrak{L}_{D}\omega^{\mu}_{h}=\omega_{h}^{\mu}\quad\textrm{for all}\;\mu=0,\cdots\,,d-1$
As in the Lagrangian case, we refer to $\Gamma\in\mathbb{R}$ as the degree of the symmetry, and employ the notation $D$ and $\Gamma$, so as to make a distinction from the corresponding Lagrangian objects $\Sigma$ and $\Lambda$.
The Hamiltonian reduction may, in principle, be carried out without reference to the Lagrangian function; however, in practice, it is more convenient to work with the Hamiltonian obtained as a result of a Legendre transform of $\widehat{L}$, introduced in (52). Writing
$\widehat{L}=e^{\rho}f(\rho_{\mu},\psi,\psi_{\mu})$
we see that
$\mathcal{FL}^{*}p^{\mu}_{\rho}=\frac{\partial\widehat{L}}{\partial\rho_{\mu}}=e^{\rho}\frac{\partial f}{\partial\rho_{\mu}}=e^{\rho}s^{\mu}\quad\quad\quad\quad\mathcal{FL}^{*}p^{\mu}_{\psi}=\frac{\partial\widehat{L}}{\partial\psi_{\mu}}=e^{\rho}\frac{\partial f}{\partial\psi_{\mu}}$ (56)
Recall that in the coordinates $(\rho,\rho_{\mu},\psi,\psi_{\mu})$, the Lagrangian scaling symmetry adopted the particularly simple form $\Sigma=\partial_{\rho}$; by virtue of this structure, we find that the corresponding vector field $D$ (with $\Gamma=1$) for the Hamiltonian is
$D=\frac{\partial}{\partial\rho}+p^{\mu}_{\rho}\frac{\partial}{\partial p^{\mu}_{\rho}}+p^{\mu}_{\psi}\frac{\partial}{\partial p^{\mu}_{\psi}}$ (57)
For the case of multiple unscaled fields $\psi^{a}$ ($a=1,\,\cdots,k$), this vector field generalises trivially as the sum of logarithmic derivatives of the momenta.
In order to obtain the symmetry-reduced Hamiltonian, we introduce (omitting the pullback by the appropriate submersion)
$s^{\mu}=\frac{p^{\mu}_{\rho}}{e^{\rho}}\quad\quad\quad\Pi^{\mu}_{\psi}=\frac{p^{\mu}_{\psi}}{e^{\rho}}$ (58)
Finally, we claim that, upon defining [^7] Here, we have made explicit the dependence of $H^{c}$ and $\eta$ on $\psi$, $\Pi^{\mu}_{\psi}$, and $s^{\mu}$ to emphasise that these are to be considered the coordinates on the reduced space.
$H^{c}(\psi,\Pi^{\mu}_{\psi},s^{\mu})=\frac{H}{e^{\rho}}\quad\quad\quad\quad\eta^{\mu}(\psi,\Pi^{\nu}_{\psi},s^{\nu})=\frac{\iota_{D}\omega^{\mu}_{h}}{e^{\rho}}=ds^{\mu}-\Pi^{\mu}_{\psi}\,d\psi$ (59)
we obtain a (pre-)multicontact structure, whose corresponding Hamiltonian $d$-form $\Theta_{H^{c}}$ is given by
$\Theta_{H^{c}}=\eta^{\mu}\wedge\,d^{d-1}x_{\mu}+H^{c}d^{d}x$ (60)
Moreover, as we shall show, the Herglotz De-Donder Weyl equations for holonomic sections correctly reproduce the dynamics derived from the full multisymplectic system, and further, they do so without reference to the scaling variable $\rho$. For simplicity, we shall continue to work with the single field variable $\psi$, together with its corresponding momenta $\Pi^{\mu}_{\psi}$; the generalisation to multiple fields requires the trivial replacement $\psi\rightarrow\psi^{a}$ and $\Pi^{\mu}_{\psi}\rightarrow\Pi^{\mu}_{a}$, with $a=1,\,\cdots,k$.
Suppose that $\gamma:M\rightarrow J^{1}\widetilde{E}^{*}\times\mathbb{R}^{d}$ is a critical section of the multicontact variational problem, that we write in local coordinates as $\gamma(x)=\left(x^{\mu},\psi(x),\Pi^{\mu}_{\psi}(x),s^{\mu}(x)\right)$; from (39), we have
$0$ $=\frac{\partial\Pi^{\mu}_{\psi}}{\partial x^{\mu}}+\left(\frac{\partial H^{c}}{\partial\psi}+\Pi^{\mu}_{\psi}\frac{\partial H^{c}}{\partial s^{\mu}}\right)$
$=\frac{\partial}{\partial x^{\mu}}\left(e^{-\rho}p^{\mu}_{\psi}\right)+\left(e^{-\rho}\frac{\partial H}{\partial\psi}+e^{-\rho}p^{\mu}_{\psi}\frac{\partial H^{c}}{\partial s^{\mu}}\right)$
$=e^{-\rho}\left(-\rho_{\mu}p^{\mu}_{\psi}+\frac{\partial p^{\mu}_{\psi}}{\partial x^{\mu}}+\frac{\partial H}{\partial\psi}+p^{\mu}_{\psi}\frac{\partial H^{c}}{\partial s^{\mu}}\right)$
$=e^{-\rho}\left(\frac{\partial p^{\mu}_{\psi}}{\partial x^{\mu}}+\frac{\partial H}{\partial\psi}\right)$
In passing to the final line, we have used (48), noting that if $\rho_{\mu}=-\frac{\partial L^{H}}{\partial s^{\mu}}$ then $\rho_{\mu}=\frac{\partial H^{c}}{\partial s^{\mu}}$.
It is thus clear that the Herglotz De-Donder Weyl equation for $\Pi^{\mu}_{\psi}$ correctly implies the corresponding multisymplectic expression for the momenta $p^{\mu}_{\psi}$. The proof for the $\psi$ equation of motion proceeds analogously, and is thus left as an exercise. The dynamical evolution of the action density, which is given by
$\frac{\partial s^{\mu}}{\partial x^{\mu}}-\left(\Pi^{\mu}_{\psi}\frac{\partial H^{c}}{\partial\Pi^{\mu}_{\psi}}+H^{c}\right)=0$
is precisely the statement that $\partial_{\mu}s^{\mu}=L^{H}$, expressed in canonically-conjugate variables.
From our construction, we have identified a symmetry of the covariant configuration space $E$, which was lifted to a vector field on the first jet bundle $J^{1}E$. Upon adopting coordinates along the scaling direction, and performing a Legendre transform, we observed that the corresponding Hamiltonian scaling symmetry $D$ included, in addition to the pure scaling $\partial_{\rho}$, further logarithmic derivative terms for each of the momentum directions. Geometrically, the reduction process appears to take place on the multiphase space, since we make use of the Hamiltonian function $H$ and the 2-forms $\omega^{\mu}_{h}$, both of which are defined on $J^{1}E^{*}$. However, the effect of the elimination of the scaling variable $\rho$ is to produce a covariant configuration space $\widetilde{E}=E/\mathbb{R}_{+}$ of dimension $\textrm{dim}\,E-1$. The reduced Hamiltonian $H^{c}$ and 1-forms $\eta^{\mu}$ are not defined on the first jet bundle of this new configuration space, but rather a manifold whose topology is of the form $J^{1}\widetilde{E}^{*}\times\mathbb{R}^{d}$.
This occurs precisely because the $d$ momenta $p^{\mu}_{\rho}$, which are no longer canonically conjugate to any field variable, assume the role of the action density, spanning a Reeb distribution [65], thereby producing a reduced space which is a multicontact manifold. Within the Lagrangian formalism, we observe a similar process; in this case, rather than the momenta $p^{\mu}_{\rho}$, it is the $d$ coordinate derivatives $\rho_{\mu}$ which span the Reeb distribution, yielding a manifold of the form $J^{1}\widetilde{E}\times\mathbb{R}^{d}$.
V Field Theory Classification
In order to better understand the geometric insight offered by our symmetry reduction procedure, it is instructive to consider the different multisymplectic bundle structures that arise within a variety of physically interesting field theories. Since the vast majority of such models are gauge theories [86, 87], we shall attempt to construct a broad system of classification, based upon the gauge group, and thence deduce important features common to each type. Our discussion will closely follow that presented in [88] and [89]; the notation employed throughout remains consistent with that developed over the course of sections (II.2) and (II.3).
In general, we shall consider two principal classes of theory, whose distinction is determined by the relationship between the gauge group $G\subset\textrm{Aut}(E)$, and the spacetime diffeomorphism group $\textrm{Diff}(M)$. Given a diffeomorphism $f:M\rightarrow M$, we say that an automorphism $\Phi:E\rightarrow E$ covers $f$ if $\pi\circ\Phi=f\circ\pi$. Intuitively, the automorphism can be considered to ‘lift’ the diffeomorphism from the spacetime manifold to the configuration space, in a manner which is consistent with the action of $f$ on $M$ [90].
A theory is described as being parameterised if all (or a suitably large subset of) spacetime diffeomorphisms are covered by the transformations in $G$. Physically, this means that the theory is invariant under arbitrary coordinate reparameterisations; this is the case of General Relativity, for example, in which the entirety of $\textrm{Diff}(M)$ is covered by $G$.
A background theory, by contrast, refers to any system in which the gauge group is a subset of those automorphisms which cover only the identity on $M$. A Yang-Mills theory defined on a flat background spacetime, for instance, has as its gauge group a set of transformations which are completely internal, and thus unrelated to spacetime diffeomorphisms.
All theories we shall consider throughout will possess a metric; however, the precise status of this metric within the particular field theory of interest will affect the underlying multisymplectic bundle structure. In the simplest of cases, the metric is a fixed object, so that it neither evolves dynamically, nor may it assume different values. Whenever we consider theories defined on a flat static background, the metric is fixed, and will most often be the Minkowski metric $\eta_{\mu\nu}$.
Another possibility is that the metric enters the theory parametrically; in this case, it is not constrained to be fixed, and we may choose to assign different values to the metric, as the context befits. Whilst this kind of theory does allow the metric to be changed, the equations of motion are derived with the metric held fixed. For this reason, we must include the components of the metric as additional configuration space variables; however, since these are not variational in nature, the Lagrangian should not contain any metric velocity variables. Consequently, when treating theories of this type, we modify the configuration space $E$ and the corresponding jet bundle $J^{1}E$ in the following way
$E\;\longrightarrow\;E\times_{M}\textrm{Sym}_{2}^{d-1,1}(M)\quad\quad\quad\quad\quad J^{1}E\;\longrightarrow\;J^{1}E\times_{M}\textrm{Sym}_{2}^{d-1,1}(M)$
in which $\textrm{Sym}_{2}^{d-1,1}(M)$ denotes the space of symmetric covariant tensors of rank two and Lorentzian signature (d1,1)(d-1,1) on $M$; sections of $\textrm{Sym}_{2}^{d-1,1}(M)$ then correspond to metrics on $M$. Note that we do not consider the jet bundle over the whole space $E\times_{M}\textrm{Sym}_{2}^{d-1,1}(M)$, but rather append the factor of $\textrm{Sym}_{2}^{d-1,1}(M)$ to the original jet bundle; this is precisely so that the components $g_{\mu\nu}$ of the metric are amongst the field variables, but the derivatives $g_{\mu\nu,\rho}$ are not.
The final possibility we shall consider is that of a variational metric; typically, theories which possess variational metrics are coupled to gravity, and are thus described by Lagrangians which do incorporate derivatives of the metric. As an example, we consider a principal $G$ bundle $\pi:P\rightarrow M$, with gauge group $G=U(1)$; denote by $p:C(P)\rightarrow M$ the bundle of connections of $P$, which we may identify with the quotient manifold $C(P)\cong J^{1}P/G$ [91]. Coupling this $U(1)$ gauge theory to gravity requires us to work with the following covariant configuration space
$E=C(P)\times_{M}\textrm{Sym}^{3,1}_{2}(M)$
upon which $(x^{\mu},A_{\mu};g_{\mu\nu})$ provide local coordinates. For concreteness, we have here restricted ourselves to a $U(1)$ gauge theory in $d=3+1$ spacetime dimensions. The gauge-invariant Lagrangian is now taken to be a function on
$\mathcal{P}:=J^{1}C(P)\times_{M}J^{1}(\textrm{Sym}_{2}^{3,1}(M))$
Note that this space differs from that which we use when describing theories whose metrics are parametric; the use of the jet bundle $J^{1}(\textrm{Sym}_{2}^{3,1}(M))$ introduces velocity variables for the components $g_{\mu\nu}$, with the result that local coordinates on $\mathcal{P}$ are given by $(x^{\mu},A_{\mu},g_{\mu\nu},A_{\mu,\nu},g_{\mu\nu,\rho})$; when evaluated on the jet prolongation $j^{1}\phi:M\rightarrow\mathcal{P}$ of some section $\phi\in\Gamma(M,E)$, the velocity coordinates become the familiar derivatives $\partial_{\nu}A_{\mu}$ and $\partial_{\rho}g_{\mu\nu}$.
Example I: N Real Scalar Fields
As an illustration of the formalism we have developed thus far, let us consider the example of $N$ non-interacting real scalar fields $\phi^{a}$ (with $a=1,\,\cdots,N$) moving in some curved, $d$-dimensional spacetime $M$, equipped with a Lorentzian metric $g_{\mu\nu}$ of signature [^8] Our use of the mostly-minus convention is primarily so that we may retain a positive kinetic term; when studying General Relativity, we shall revert to the more customary (,+,+,)(-$\cdots$) signature. $(+,-,-,\cdots)$. Whilst we allow for a background different from Minkowski spacetime, we shall impose that the metric be non-variational, and thus parametric in nature. In accordance with the discussion of the previous section, this requires us to make the replacements
$E\;\rightarrow\;\mathcal{E}:=E\times_{M}\textrm{Sym}^{d-1,1}_{2}(M)\quad\quad\quad\quad J^{1}E\;\rightarrow\;\mathcal{C}:=J^{1}E\times_{M}\textrm{Sym}^{d-1,1}_{2}(M)$
where $E$ is the trivial bundle $E=M\times\mathbb{R}^{N}\rightarrow M$, which would be an appropriate covariant configuration space for a flat background. Consequently, local coordinates on $\mathcal{C}$ are $(x^{\mu},\phi^{a},\phi^{a}_{\mu};g_{\mu\nu})$, where we have used a semicolon to separate the variational and parametric degrees of freedom. The Lagrangian function $L:\mathcal{C}\rightarrow\mathbb{R}$ is then given by
$L=\sqrt{-g}\sum_{a=1}^{N}\left[\frac{1}{2}g^{\mu\nu}\phi_{\mu}^{a}\phi_{\nu}^{a}-\frac{1}{2}m^{2}(\phi^{a})^{2}\right]$ (61)
Note that the $N$ scalar fields $\phi^{a}$ may be collected into an $N$-tuplet, making manifest the global $O(N)$ symmetry of (61). For the purpose of facilitating the reduction process, we shall make a change of coordinates, such that this $O(N)$ is no longer manifest; in particular, we shall adopt $N$-dimensional polar coordinates $(R,X^{1},\,\cdots\,,X^{N-1})$, defined according to
$\begin{split}\phi^{1}=R\cos X^{1}\\ \phi^{2}=R\sin X^{1}\cos X^{2}\\ \vdots\quad\quad\vdots\quad\quad\vdots\quad\quad\\ \phi^{N-1}=R\sin X^{1}\cdots\,\sin X^{N-2}\cos X^{N-1}\\ \phi^{N}=R\sin X^{1}\cdots\,\sin X^{N-2}\sin X^{N-1}\end{split}$ (62)
Writing the radial coordinate $R$ as $e^{\rho/2}$, the Lagrangian (61) may be expressed as
$L=e^{\rho}\sqrt{-g}\left[g^{\mu\nu}\left(\frac{1}{8}\rho_{\mu}\rho_{\nu}+\sum_{a=1}^{N-1}\frac{1}{2}X_{\mu}^{a}X_{\nu}^{a}\right)-\frac{1}{2}m^{2}\right]$ (63)
The Poincaré-Cartan form $\Theta_{L}$ is simply
$\Theta_{L}=e^{\rho}\sqrt{-g}\,g^{\mu\nu}\left(\frac{1}{4}\rho_{\nu}\,d\rho+\sum_{a=1}^{N-1}X^{a}_{\nu}\,dX^{a}\right)\wedge\,d^{3}x_{\mu}-e^{\rho}\sqrt{-g}\left[g^{\mu\nu}\left(\frac{1}{8}\rho_{\mu}\rho_{\nu}+\sum_{a=1}^{N-1}\frac{1}{2}X_{\mu}^{a}X_{\nu}^{a}\right)+\frac{1}{2}m^{2}\right]d^{4}x$
from which, in accordance with the procedure outlined in section (IV.1), we compute the 1-forms $\theta_{L}^{\mu}$
$\theta_{L}^{\mu}=e^{\rho}\sqrt{-g}\,g^{\mu\nu}\left(\frac{1}{4}\rho_{\nu}\,d\rho+\sum_{a=1}^{N-1}X^{a}_{\nu}\,dX^{a}\right)$
It is then apparent that the vector field $\Sigma=\partial_{\rho}$ satisfies $\mathfrak{L}_{\Sigma}L=L$ and $\mathfrak{L}_{\Sigma}\theta_{L}^{\mu}=\theta_{L}^{\mu}$, and so is a scaling symmetry of degree one. The function $f$ used to construct the Herglotz Lagrangian in (53) is trivially read off from (63), and we deduce the action density to be
$s^{\mu}=\frac{\partial f}{\partial\rho_{\mu}}=\frac{1}{4}\sqrt{-g}\,g^{\mu\nu}\rho_{\nu}$
Finally, we eliminate the derivative coordinate $\rho_{\mu}$ in favour of the action density, in order to obtain the following Herglotz Lagrangian
$L^{H}=\sqrt{-g}\left[\frac{1}{2}\sum_{a=1}^{N-1}g^{\mu\nu}X^{a}_{\mu}X^{a}_{\nu}-\frac{1}{2}m^{2}\right]-\frac{2}{\sqrt{-g}}g_{\mu\nu}s^{\mu}s^{\nu}$ (64)
Geometrically speaking, this Herglotz Lagrangian is a function on the space
$\mathcal{S}=(J^{1}\widetilde{E}\times_{M}\textrm{Sym}^{d-1,1}_{2}(M))\times\mathbb{R}^{d}$ (65)
in which $\widetilde{E}\cong M\times\mathbb{R}^{N-1}$ is a configuration space formed by taking the quotient of $E=M\times\mathbb{R}^{N}$ by the one-dimensional orbits of $\Sigma=\partial_{\rho}$. Additionally, $L^{H}$ is insensitive to internal rotations amongst the $X^{a}$, and so the reduced theory possesses a global O(N1)O(N-1) symmetry.
We have written the Herglotz Lagrangian in a way which makes manifest the elegant physical interpretation of our reduction procedure. We began with a conservative theory of $N$ massive, non-interacting scalar fields; this system possessed a scaling symmetry, in which the superfluous degree of freedom could be considered to be a global radial scale. Upon eliminating this variable, we obtain a theory of (N1)(N-1) massless scalar fields subject to a constant potential, whose strength is fixed by the original mass, together with a friction-like term which arises due to the inaccessibility of the radial scale. Moreover, from the form of the Lagrangian, it is apparent that the massless scalar dynamics appear completely decoupled from the friction-like piece.
One may be inclined to object that the two components are in fact coupled, due to the presence of the factor of $\textrm{Sym}_{2}^{d-1,1}(M)$. However, this not the case, as the background is parametric, and so $\textrm{Sym}_{2}^{d-1,1}(M)$ simply describes the space of all possible choices of metric, and is in no way dynamical. In light of this, we may write the Herglotz Lagrangian (64) as the sum of two independent pieces: $L^{H}=L^{H}_{N-1}+L^{H}_{\textrm{fric}}$, in which $L^{H}_{N-1}$ is the Lagrangian of the (N1)(N-1) massless scalar fields on $J^{1}\widetilde{E}\times_{M}\textrm{Sym}^{d-1,1}_{2}(M)$, and $L^{H}_{\textrm{fric}}$ refers to the frictional component, parameterised by the action density.
This separation into a Lagrangian on a lower-dimensional jet bundle and an action-dependent piece is highly attractive from the perspective of quantisation. Unfortunately, it is far from clear how such a process should proceed, and, as we shall see in the second example, the decoupling of the frictional component from the reduced space dynamics does not survive within an interacting field theory.
Let us now examine the corresponding contact Hamiltonian; the most straightforward way to obtain $H^{c}$ would be to perform a Legendre transform on (64); however, for didactical purposes, we shall compute the full Hamiltonian from (63), and carry out the reduction procedure using the methods of section (IV.2).
The field momenta $p^{\mu}_{\rho}$ and $p^{\mu}_{a}$ corresponding to $\rho$ and $X^{a}$ respectively are
$p^{\mu}_{\rho}=\frac{\partial L}{\partial\rho_{\mu}}=\frac{1}{4}e^{\rho}\sqrt{-g}\,g^{\mu\nu}\rho_{\nu}\quad\quad\quad\quad\quad p^{\mu}_{a}=\frac{\partial L}{\partial X^{a}_{\mu}}=e^{\rho}\sqrt{-g}\sum_{a=1}^{N-1}g^{\mu\nu}X_{\nu}^{a}$ (66)
Accordingly, the Hamiltonian function corresponding to (63) is given by
$H=\frac{1}{e^{\rho}\sqrt{-g}}g_{\mu\nu}\left(2p^{\mu}_{\rho}p^{\nu}_{\rho}+\frac{1}{2}\sum_{a=1}^{N-1}p_{a}^{\mu}p_{a}^{\nu}\right)+\frac{1}{2}e^{\rho}\sqrt{-g}\,m^{2}$ (67)
From (15), we have the 2-forms
$\omega_{h}^{\mu}=d\rho\wedge dp^{\mu}_{\rho}+\sum_{a=1}^{N-1}dX^{a}\wedge dp_{a}^{\mu}$
and, as expected, the vector field
$D=\frac{\partial}{\partial\rho}+p^{\mu}_{\rho}\frac{\partial}{\partial p^{\mu}_{\rho}}+\sum_{a=1}^{N-1}p^{\mu}_{a}\frac{\partial}{\partial p^{\mu}_{a}}$ (68)
is a scaling symmetry of degree one. The contact Hamiltonian and 1-forms $\eta^{\mu}$ are easily obtained, upon introducing the following coordinates on the reduced space
$s^{\mu}=\frac{p^{\mu}_{\rho}}{e^{\rho}}\quad\quad\quad\quad\quad\Pi^{\mu}_{a}=\frac{p^{\mu}_{a}}{e^{\rho}}$ (69)
We then find that
$H^{c}=\frac{H}{e^{\rho}}=\frac{1}{2\sqrt{-g}}\sum_{a=1}^{N-1}g_{\mu\nu}\Pi^{\mu}_{a}\Pi^{\nu}_{a}+\frac{1}{2}\sqrt{-g}\,m^{2}+\frac{2}{\sqrt{-g}}g_{\mu\nu}s^{\mu}s^{\nu}\quad\quad\quad\quad\eta^{\mu}=\frac{\iota_{D}\omega^{\mu}_{h}}{e^{\rho}}=ds^{\mu}-\sum_{a=1}^{N-1}\Pi^{\mu}_{a}\,dX^{a}$ (70)
The multiphase space upon which the original Hamiltonian (67) was defined is $J^{1}E^{*}\times_{M}\textrm{Sym}_{2}^{d-1,1}(M)$, in which we recall that $J^{1}E^{*}$ is the multimomentum bundle over $E=M\times\mathbb{R}^{N}$. Having made the contact reduction, $H^{c}$ is now a function on
$\mathcal{S}^{*}=(J^{1}\widetilde{E}^{*}\times_{M}\textrm{Sym}_{2}^{d-1,1}(M))\times\mathbb{R}^{d}$ (71)
where, as in (65), $\widetilde{E}=M\times\mathbb{R}^{N-1}$. Unsurprisingly, as in the Lagrangian formalism, we observe that our contact Hamiltonian (70) consists of a piece describing a theory of (N1)(N-1) massless scalar fields, with a constant shift in the energy, together with an action-dependent term, whose dissipative nature serves to compensate the inaccessibility of the global scale variable.
Example II: Interacting Scalar Fields
In the previous example, we considered the case of $N$ non-interacting massive scalar fields, and observed that, dynamically, such a system is entirely equivalent to a configuration of (N1)(N-1) non-interacting massless scalar fields, together with a decoupled friction-like component. Any field theory of physical interest necessarily contains interactions amongst its components; for this reason, we shall now analyse a simple model of two real, interacting scalar fields $\phi$ and $\chi$.
Working over the two-dimensional spacetime manifold $M=\mathbb{R}^{1,1}$, with coordinates $(t,x)$, and fixed Minkowski metric $\eta_{\mu\nu}=\textrm{diag}(1,-1)$, we have the covariant configuration space $E=M\times\mathbb{R}^{2}$. The first jet bundle $J^{1}E$ admits local coordinates $(t,x,\phi,\chi,\phi_{\mu},\chi_{\mu})$, and the Lagrangian function we shall consider is
$L=\frac{1}{2}\left(\phi_{t}^{2}-\phi^{2}_{x}\right)+\frac{1}{2}\phi^{2}\left(\chi_{t}^{2}-\chi^{2}_{x}\right)+\phi\,(\phi_{t}-\phi_{x})(\chi_{t}-\chi_{x})-\chi\phi^{2}$ (72)
Here, since we have only two spacetime coordinates, we have chosen to label the velocities as $\phi_{t}$, rather than $\phi_{0}$, for example. Evidently, this Lagrangian is highly non-physical, and should therefore be considered to be nothing more than a convenient and illustrative example. Additionally, there are a number of dimensionful coupling constants, which, for simplicity of exposition, have been set to unity. In the previous example, significant emphasis was placed on the geometry of the symmetry-reduced space, and subsequent physical interpretation of the final theory; in order to ensure an adequate presentation of all facets of the formalism we have developed throughout, we shall here strive to illustrate the process by which multivector field solutions to the equations of motion are obtained.
It is relatively easy to see from (72) that $\phi$ is our scaling variable; indeed, in keeping with the notation developed throughout section (IV.1), we have
$\bar{\Sigma}=\frac{1}{2}\phi\frac{\partial}{\partial\phi}\quad\quad\implies\quad\quad\Sigma=\frac{1}{2}\left(\phi\frac{\partial}{\partial\phi}+\phi_{t}\frac{\partial}{\partial\phi_{t}}+\phi_{x}\frac{\partial}{\partial\phi_{x}}\right)$
which is a scaling symmetry of degree one. Adopting coordinates along the scaling direction, we change field variables according to $\phi\mapsto e^{\rho/2}$. Consequently, the Lagrangian (72) may be expressed as
$L=e^{\rho}\left[\frac{1}{8}(\rho_{t}^{2}-\rho_{x}^{2})+\frac{1}{2}(\chi_{t}^{2}-\chi_{x}^{2})+\frac{1}{2}(\rho_{t}-\rho_{x})(\chi_{t}-\chi_{x})-\chi\right]$ (73)
A short calculation shows that the Hessian matrix for this system is given by
$W=\frac{e^{\rho}}{4}\begin{pmatrix}1&0&2&-2\\ 0&-1&-2&2\\ 2&-2&4&0\\ -2&2&0&-4\end{pmatrix}$
which is of constant rank four, confirming that $L$ is a regular Lagrangian. The corresponding multisymplectic 2-form $\Theta_{L}$ is given by
$\begin{split}\Theta_{L}=e^{\rho}\biggr[\left(\frac{1}{4}\rho_{t}+\frac{1}{2}\left(\chi_{t}-\chi_{x}\right)\right)\,d\rho\wedge dx+\left(\frac{1}{4}\rho_{x}+\frac{1}{2}\left(\chi_{t}-\chi_{x}\right)\right)\,d\rho\wedge dt+\left(\chi_{t}+\frac{1}{2}\left(\rho_{t}-\rho_{x}\right)\right)\,d\chi\wedge dx\\ +\,\left(\chi_{x}+\frac{1}{2}\left(\rho_{t}-\rho_{x}\right)\right)\,d\chi\wedge dt-\left(\frac{1}{8}(\rho_{t}^{2}-\rho_{x}^{2})+\frac{1}{2}(\chi_{t}^{2}-\chi_{x}^{2})+\frac{1}{2}(\rho_{t}-\rho_{x})(\chi_{t}-\chi_{x})+\chi\right)dt\wedge dx\biggr]\end{split}$ (74)
We shall now conduct an analysis of the dynamics described by this multisymplectic Lagrangian system, before carrying out the symmetry reduction. We then conclude this example, comparing the on-shell behaviour of each of the two theories. A straightforward computation of the Euler-Lagrange field equations for $\rho$ yields
$\frac{1}{8}\hskip-1.42262pt\left[(\partial_{t}\rho)^{2}-(\partial_{x}\rho)^{2}\right]-\frac{1}{2}\hskip-1.42262pt\left[(\partial_{t}\chi)^{2}-(\partial_{x}\chi)^{2}\right]+\frac{1}{4}(\partial_{t}^{2}-\partial_{x}^{2})\rho+\frac{1}{2}(\partial_{t}-\partial_{x})^{2}\chi+\chi=0$ (75)
Similarly, for $\chi$, we find that
$\frac{1}{2}\left(\partial_{t}\rho-\partial_{x}\rho\right)^{2}+\frac{1}{2}(\partial_{t}-\partial_{x})^{2}\rho+\partial_{t}\rho\,\partial_{t}\chi-\partial_{x}\rho\,\partial_{x}\chi+(\partial_{t}^{2}-\partial_{x}^{2})\chi+1=0$ (76)
In order to illustrate how these equations of motion may be deduced from a more geometrical perspective, we shall assume that there exist locally decomposable $\widehat{\pi}-$transverse multivector field solutions $\boldsymbol{X}_{L}\in\textrm{ker}_{\widehat{\pi}}^{2}\,\Omega_{L}$, and write
$\begin{split}\boldsymbol{X}_{L}=\left(\frac{\partial}{\partial t}+F^{\rho}_{t}\frac{\partial}{\partial\rho}+F^{\chi}_{t}\frac{\partial}{\partial\chi}+G^{\rho}_{tt}\frac{\partial}{\partial\rho_{t}}+G^{\rho}_{tx}\frac{\partial}{\partial\rho_{x}}+G^{\chi}_{tt}\frac{\partial}{\partial\chi_{t}}+G^{\chi}_{tx}\frac{\partial}{\partial\chi_{x}}\right)\;\wedge\\ \left(\frac{\partial}{\partial x}+F^{\rho}_{x}\frac{\partial}{\partial\rho}+F^{\chi}_{x}\frac{\partial}{\partial\chi}+G^{\rho}_{xt}\frac{\partial}{\partial\rho_{t}}+G^{\rho}_{xx}\frac{\partial}{\partial\rho_{x}}+G^{\chi}_{xt}\frac{\partial}{\partial\chi_{t}}+G^{\chi}_{xx}\frac{\partial}{\partial\chi_{x}}\right)\end{split}$ (77)
Here, $\widehat{\pi}-$transversality has been enforced through the condition $\iota_{\scriptscriptstyle\boldsymbol{X}_{L}}(dt\wedge dx)=1$, which sets the multiplicative function $f$ in (6) to unity. Substituting this local decomposition into the equation $\iota_{\scriptscriptstyle\boldsymbol{X}_{L}}\Omega_{L}=0$, we obtain
$0$ $=\frac{1}{8}(\rho_{t}^{2}-\rho_{x}^{2})+\frac{1}{2}(\chi_{t}^{2}-\chi_{x}^{2})+\frac{1}{2}(\rho_{t}-\rho_{x})(\chi_{t}-\chi_{x})+\chi-F^{\chi}_{t}\left(\chi_{t}+\frac{1}{2}\left(\rho_{t}-\rho_{x}\right)\right)$ (78a)
$+F^{\chi}_{x}\left(\chi_{x}+\frac{1}{2}\left(\rho_{t}-\rho_{x}\right)\right)+\frac{1}{4}\left(G^{\rho}_{tt}-G^{\rho}_{xx}\right)+\frac{1}{2}\left(G^{\chi}_{tt}-G^{\chi}_{tx}-G^{\chi}_{xt}+G^{\chi}_{xx}\right)$
$0$ $=F^{\rho}_{t}\left(\chi_{t}+\frac{1}{2}\left(\rho_{t}-\rho_{x}\right)\right)-F^{\rho}_{x}\left(\chi_{x}+\frac{1}{2}\left(\rho_{t}-\rho_{x}\right)\right)+\frac{1}{2}\left(G^{\rho}_{tt}-G^{\rho}_{tx}-G^{\rho}_{xt}+G^{\rho}_{xx}\right)$ (78b)
$+G^{\chi}_{tt}-G^{\chi}_{xx}+1$
$0$ $=\frac{1}{4}\left(\rho_{t}-F^{\rho}_{t}\right)+\frac{1}{2}\biggr[\left(\chi_{t}-\chi_{x}\right)-\left(F^{\chi}_{t}-F^{\chi}_{x}\right)\biggr]$ (78c)
$0$ $=\frac{1}{4}\left(\rho_{x}-F^{\rho}_{x}\right)+\frac{1}{2}\biggr[\left(\chi_{t}-\chi_{x}\right)-\left(F^{\chi}_{t}-F^{\chi}_{x}\right)\biggr]$ (78d)
$0$ $=\left(\chi_{t}-F^{\chi}_{t}\right)+\frac{1}{2}\biggr[\left(\rho_{t}-\rho_{x}\right)-\left(F^{\rho}_{t}-F^{\rho}_{x}\right)\biggr]$ (78e)
$0$ $=\left(\chi_{x}-F^{\chi}_{x}\right)+\frac{1}{2}\biggr[\left(\rho_{t}-\rho_{x}\right)-\left(F^{\rho}_{t}-F^{\rho}_{x}\right)\biggr]$ (78f)
Since the Lagrangian system under consideration is regular, the self-consistency of these expressions is assured; imposing semi-holonomy of our solutions requires that we set
$\begin{cases}\;\;F^{\rho}_{t}=\rho_{t}\quad\quad\quad F^{\chi}_{t}=\chi_{t}\\ \;\;F^{\rho}_{x}=\rho_{x}\quad\quad\quad F^{\chi}_{x}=\chi_{x}\end{cases}$
Clearly equations (78c) - (78f) are rendered trivial by this condition, whilst the first two may now be expressed as
$0$ $=\frac{1}{8}(\rho_{t}^{2}-\rho_{x}^{2})-\frac{1}{2}(\chi_{t}^{2}-\chi_{x}^{2})+\chi+\frac{1}{4}\left(G^{\rho}_{tt}-G^{\rho}_{xx}\right)+\frac{1}{2}\left(G^{\chi}_{tt}-G^{\chi}_{tx}-G^{\chi}_{xt}+G^{\chi}_{xx}\right)$ (79a)
$0$ $=\chi_{t}^{2}-\chi_{x}^{2}+\frac{1}{2}(\rho_{t}-\rho_{x})(\chi_{t}-\chi_{x})+\frac{1}{2}\left(G^{\rho}_{tt}-G^{\rho}_{tx}-G^{\rho}_{xt}+G^{\rho}_{xx}\right)+G^{\chi}_{tt}-G^{\chi}_{xx}+1$ (79b)
If our multivector field $\boldsymbol{X}_{L}=X_{t}\wedge X_{x}$ is holonomic, we require that the distribution generated by $X_{t}$ and $X_{x}$ be involutive. In such a case, we have integral sections $\psi:M\rightarrow E$, which we represent locally as $\psi(t,x)=(t,x,\rho(t,x),\chi(t,x))$; then the multivector field components $G^{\rho}_{\mu\nu}$ and $G^{\chi}_{\mu\nu}$ satisfy
$G^{\rho}_{\mu\nu}=\frac{\partial^{2}\rho}{\partial x^{\mu}\partial x^{\nu}}\quad\quad\textrm{and}\quad\quad G^{\chi}_{\mu\nu}=\frac{\partial^{2}\chi}{\partial x^{\mu}\partial x^{\nu}}$
Upon making these substitutions in (79a) and (79b), we recover precisely the Euler-Lagrange field equations (75) and (76).
Having analysed the multisymplectic Lagrangian system, we may now make the symmetry reduction to obtain the Herglotz Lagrangian. From (73), we immediately read off the function $f(\rho_{\mu},\chi,\chi_{\mu})$, from which we obtain the following action density
$s^{\mu}:=\frac{\partial f}{\partial\rho_{\mu}}\quad\quad\quad\quad\quad s^{t}=\frac{1}{4}\rho_{t}+\frac{1}{2}\left(\chi_{t}-\chi_{x}\right)\quad\quad\quad s^{x}=-\left(\frac{1}{4}\rho_{x}+\frac{1}{2}\left(\chi_{t}-\chi_{x}\right)\right)$ (80)
Constructing the Herglotz Lagrangian $L^{H}=f-\rho_{\mu}s^{\mu}$, and eliminating reference to the velocity coordinates $\rho_{\mu}$, we find that
$L^{H}=\frac{1}{2}(\chi_{t}^{2}-\chi_{x}^{2})-2\left[(s^{t})^{2}-(s^{x})^{2}\right]+2(s^{t}+s^{x})(\chi_{t}-\chi_{x})-\chi$ (81)
The Herglotz-Lagrange field equations for $\chi$ then read
$\partial_{t}(\partial_{t}\chi+2(s^{t}+s^{x}))-\partial_{x}(\partial_{x}\chi+2(s^{t}+s^{x}))+1=(\partial_{t}\chi-\partial_{x}\chi)^{2}-4(s^{t}\partial_{t}\chi+s^{x}\partial_{x}\chi)-8(s^{t}+s^{x})^{2}$ (82)
It is a straightforward exercise to show that, with the expressions (80) for the components of the action density, this system correctly reproduces the dynamical evolution of the field $\chi$ as described by (75) and (76). From this Herglotz Lagrangian, we calculate the 2-form $\Theta_{L^{H}}$
$\begin{split}\Theta_{L^{H}}=\biggr[ds^{t}-\left(\chi_{t}+2\left(s^{t}+s^{x}\right)\right)d\chi\biggr]\wedge\,dx\,-\,\biggr[ds^{x}+\left(\chi_{x}+2\left(s^{t}+s^{x}\right)\right)d\chi\biggr]\wedge\,dt\\ +\left(\frac{1}{2}(\chi_{t}^{2}-\chi_{x}^{2})+2\left[(s^{t})^{2}-(s^{x})^{2}\right]+\chi\right)dt\wedge dx\end{split}$ (83)
together with the dissipation form
$\sigma_{\Theta}=-\,\frac{\partial L^{H}}{\partial s^{\mu}}\,dx^{\mu}=2\left(2s^{t}-\chi_{t}+\chi_{x}\right)dt-2\left(2s^{x}+\chi_{t}-\chi_{x}\right)dx$ (84)
The equations of motion could then be recovered geometrically, via an analysis similar to that carried out above; recall, however, that the multicontact 3-form $\Omega_{L^{H}}$ is constructed from the twisted differential operator $\Omega_{L^{H}}=\bar{d}\Theta_{L^{H}}$, as defined in (27). We shall not conduct such an analysis here, since the procedure is entirely analogous to that which we have already illustrated, and the results of the geometric formulation are known to be equivalent to the content of the Herglotz-Lagrange field equations.
The elimination of the scaling coordinate from this theory has an interesting physical interpretation. In the previous example, we considered the case of non-interacting scalar fields, where it was observed that the Herglotz Lagrangian assumed the form of a sum of a piece describing the reduced space scalar field dynamics, together with an independent component containing the action density variables. In this case, by contrast, when the original scalar fields exhibit some kind of interaction (which here, is mediated by the mixed derivative term $(\rho_{t}-\rho_{x})(\chi_{t}-\chi_{x})$) we observe that the corresponding Herglotz Lagrangian contains terms which couple $\chi$ to the action density. In this way, upon eliminating the scaling variable $\rho$, the coupling between the two original scalar fields manifests as an interaction between the remaining field and the frictional degrees of freedom.
Let us now consider the Hamiltonian formalism for this dynamical system; since our Lagrangian is regular, we know that the Legendre map is a diffeomorphism, and thus expect that the momenta be invertible, allowing us to unambiguously solve for the velocities $\rho_{\mu}$ and $\chi_{\mu}$. Indeed, we have that
$p^{\mu}_{\rho}=\frac{\partial L}{\partial\rho_{\mu}}\quad\quad\quad\quad p^{t}_{\rho}=e^{\rho}\left(\frac{1}{4}\rho_{t}+\frac{1}{2}\left(\chi_{t}-\chi_{x}\right)\right)\quad\quad p^{x}_{\rho}=-\,e^{\rho}\left(\frac{1}{4}\rho_{x}+\frac{1}{2}\left(\chi_{t}-\chi_{x}\right)\right)$ (85)
and similarly for $\chi$
$p^{\mu}_{\chi}=\frac{\partial L}{\partial\chi_{\mu}}\quad\quad\quad\quad p^{t}_{\chi}=e^{\rho}\left(\chi_{t}+\frac{1}{2}\left(\rho_{t}-\rho_{x}\right)\right)\quad\quad p^{x}_{\chi}=-\,e^{\rho}\left(\chi_{x}+\frac{1}{2}\left(\rho_{t}-\rho_{x}\right)\right)$ (86)
We then find that the Hamiltonian is given by
$H=2e^{-\rho}\left[(p_{\rho}^{t})^{2}-(p^{x}_{\rho})^{2}+\frac{1}{4}\left((p_{\chi}^{t})^{2}-(p^{x}_{\chi})^{2}\right)-(p^{t}_{\rho}+p^{x}_{\rho})(p^{t}_{\chi}+p^{x}_{\chi})\right]+e^{\rho}\chi$ (87)
As expected, the vector field
$D=\frac{\partial}{\partial\rho}+p^{t}_{\rho}\frac{\partial}{\partial p^{t}_{\rho}}+p^{x}_{\rho}\frac{\partial}{\partial p^{x}_{\rho}}+p^{t}_{\chi}\frac{\partial}{\partial p^{t}_{\chi}}+p^{x}_{\chi}\frac{\partial}{\partial p^{x}_{\chi}}$ (88)
constitutes a scaling symmetry of degree one; introducing the reduced space coordinates
$s^{\mu}=\frac{p^{\mu}_{\rho}}{e^{\rho}}\quad\quad\quad\quad\Pi^{\mu}_{\chi}=\frac{p^{\mu}_{\chi}}{e^{\rho}}$
we see that the contact Hamiltonian $H^{c}$ may be expressed as
$H^{c}=\frac{H}{e^{\rho}}=\frac{1}{2}\left[(\Pi^{t}_{\chi})^{2}-(\Pi^{x}_{\chi})^{2}\right]+\chi+2\left[(s^{t})^{2}-(s^{x})^{2}\right]-2(s^{t}+s^{x})(\Pi^{t}_{\chi}+\Pi^{x}_{\chi})$ (89)
A straightforward calculation shows that this contact Hamiltonian is precisely the function obtained from a direct application of the Legendre map to (81). The equations of motion derived from $H^{c}$ read
$\partial_{t}\chi$ $=\Pi^{t}_{\chi}-2(s^{t}-s^{x})$ $\partial_{x}\chi$ $=-\left(\Pi_{\chi}^{x}+2\left(s^{t}-s^{x}\right)\right)$
$\partial_{t}\Pi^{t}_{\chi}+\partial_{x}\Pi^{x}_{\chi}=2\left(\Pi^{t}_{\chi}+\Pi^{x}_{\chi}\right)^{2}-4\left(\Pi^{t}_{\chi}s^{t}-\Pi^{x}_{\chi}s^{x}\right)-1$ (90)
$\partial_{t}s^{t}+\partial_{x}s^{x}=\frac{1}{2}\left[(\Pi^{t}_{\chi})^{2}-(\Pi^{x}_{\chi})^{2}\right]-\chi-2\left[(s^{t})^{2}-(s^{x})^{2}\right]$
Note that, as must be the case, the final expression for $\partial_{\mu}s^{\mu}$ yields the Herglotz Lagrangian (81), expressed in momentum variables.
VI Conclusions and Outlook
The use of multisymplectic geometry to provide a manifestly covariant description of classical field theory has recently assumed a more prominent role within the literature, due in part to the success of standard symplectic geometry, which continues to provide insightful connections between classical mechanics and differential geometry [92]. The finite dimensionality of the multisymplectic phase space is highly attractive, and constitutes a notable advantage over the canonical field-theoretic framework. The formalism of multicontact geometry, by contrast, is far less developed; action-dependent field theories provide a way to describe systems which exhibit non-conservative effects, and are thus encountered somewhat less frequently.
Following a brief review of the mathematical preliminaries required to treat classical field theories in a manifestly covariant manner, we introduced the notion of dynamical similarity, and demonstrated that the process by which a redundant ‘scaling’ degree of freedom may be eliminated has a well-defined generalisation to a field-theoretic context. In particular, we showed that a multisymplectic Lagrangian (or Hamiltonian) system admitting a dynamical similarity could be reduced to an action-dependent theory, defined on a lower-dimensional space. From a geometrical perspective, this symmetry reduction process had a simple yet elegant interpretation: the removal of a single degree of freedom $\rho$ leaves each of the velocities $\rho_{\mu}$ with no corresponding field variable; in the reduced theory, each of these unpaired velocities assumes the role of a component of the action density $s^{\mu}$, which, together, span a Reeb distribution, making the resulting theory multicontact.
Through use of a pair of simple examples, we found that, in general, a theory of $N$ non-interacting massive scalar fields, which possesses a scaling symmetry, is dynamically-equivalent to a configuration of (N1)(N-1) massless fields moving in a non-zero potential, set by the value of the original field mass, together with a completely independent frictional component, which we interpret as arising due to the empirical inaccessibility of the scaling variable. Furthermore, it was observed that any coupling between the scalar fields of the original multisymplectic theory destroys this separation into independent pieces, and introduces terms which mix the frictional degrees of freedom with those of the remaining scalar fields.
The results presented throughout constitute a foundational framework, within which a number of interesting areas of future work may be explored. For example, we have focused exclusively on regular Lagrangian systems; the regularity of our field theory guarantees that symmetry reduced space admits a well-defined multicontact structure. The vast majority of theories of physical interest possess gauge symmetries, which give rise to singular Lagrangian systems. In this case, we must develop a formalism adapted to the treatment of such theories, and, in particular, it will be of great interest to study the interaction between gauge fixing and contact reduction. At present, it remains unclear as to whether the process of pre-symplectic reduction, used to define the physical phase space, commutes with the elimination of a scaling degree of freedom [93]. Such a framework will also provide an arena within which to study General Relativity; it has been mentioned several times that, upon making a conformal decomposition, the Einstein-Hilbert Lagrangian possesses a dynamical similarity. Having developed a sound understanding of the interaction between gauge and scaling symmetries, it will be of great interest to examine the solution space of contact-reduced General Relativity, particularly at points where our conventional description becomes singular and thus ill-defined. It is known, for instance, that within the cosmological sector, both FLRW and certain classes of Bianchi models may be extended beyond the initial singularity [94, 95].
In addition to the treatment of singular field theories, it will also be of interest to consider the behaviour of systems possessing fermionic degrees of freedom. Some progress has been made in this area; for example a De-Donder Weyl Hamiltonian description of a simple two-dimensional supersymmetric theory has been developed in [96]. Moreover, generalised Kähler geometries have been used to construct covariant Hamiltonian formulations of non-linear sigma models [97], and so an extension of our formalism to encompass anticommuting variables is highly desirable.
Finally, we have mentioned several times throughout the advantages presented by the covariant De-Donder Weyl formulation of classical field theories. Whilst manifest covariance and finite-dimensional phase spaces are indeed attractive features of this approach, when contrasted with the canonical formalism, it cannot be overlooked that the latter provides a clear and rigorous path to quantisation, whereas the former does not. In light of this, of particular importance is the development of a standardised framework within which the restricted multimomentum bundle $J^{1}E^{*}$ may be quantised, in much the same way that geometric quantisation offers a scheme for performing such calculations on the tangent bundle.